### v4

# Two axes in non-associative algebras with a Frobenius form

Abstract.
Throughout this paper is a commutative non-associative algebra over a field of characteristic not In addition posses a Frobenius form. We obtain detailed information about the multiplication in given two axes of type half in
Key words and phrases:
idempotent, axis, primitive, Jordan type, axial algebra, derivation2020 Mathematics Subject Classification:
Primary: 17C27, 17A99, Secondary: 17C101. Introduction
Throughout this paper is a commutative non-associative algebra over a field of characteristic not Assume that is a Frobenius form on that is, is a symmetric bilinear map such that We denote by the radical of . Recall that for we denote We direct the reader to Definition 1.1 below and the beginning of Β§2 for definitions required for the introduction.
This paper was inspired by [D], which in turn was inspired by [GSS]. These two papers reveal an unexpected, and rather amazing structure in primitive axial algebras of Jordan type half. Since axial algebras are related to -transposition groups, on the one hand, and to the Griess Monster algebra, on the other hand, there must be something real there (:)).
Now when one looks carefully on the arguments in [D], one realizes that all that is being used for an idempotent is that: iff is an axis (where in characteristic one has to also require see Lemma 2.3). However, some of the arguments in [D] are sleek, smart (and short), but I thought there should be more elementary (albeit longer) ones. Thus, I decided to embark on the tedious (though rewarding) effort to make the arguments elementary, and, on the way, there are some very interesting identities that are obtained.
Definitions and notation 1.1.
Recall that an idempotent is called a primitive axis of Jordan type if and Further, satisfies the following fusion rules: for and and For short we will say that is an axis.
Throughout the paper, are fixed axes in
- (3)
Throughout the paper, except for Β§2, when we write (but there is no element around) we just mean
Recall that for an axis the Miyamoto map where is an automorphism of order at most of
Theorem 1.2 (Theorem 3.7).
Let be axes. Then the following are equivalent
is a derivation on
for all and
for all
The following result is essentially a consequence of [D, Theorem 1.2 and Proposition 4.3].
Theorem 1.3 (Theorem 5.1).
If then
Theorem 1.4 (Theorem 4.7).
Let and then
Theorem 1.5 (Theorem 5.3).
Assume then
for all
For cases , we have the following theorems.
Theorem 1.7 (Theorem 7.1).
If then is a derivation.
Theorem 1.8.
We were not able to decide what happens when and (though we know that ). There is a good number of additional identities obtained in this paper. We left the case for future works.
2. The polynomials and
Lemma 2.1.
Let If then In particular where
Proof.
If then, by hypothesis, and Assume Then
Hence if then β
Lemma 2.2 (Serssβ Lemma).
Let be an axis. Then for all and
Proof.
Write and with and Then
so
while
Lemma 2.3.
Let be an idempotent, then
If and then
The following conditions on are equivalent:
for all
and
Assume Then
is a primitive axis of Jordan type iff for all
Proof.
(1):Β By definition, so (1) holds.
(2) Assume that for all Then , and Thus Also Hence,
Since we see that
Assume that holds. Since Let and By Lemma 2.1, Now and
(3)(a):Β We have
because and since is Frobenius.
(b):Β Assume that Let and By Lemma 2.1, Also since is Frobenius, Hence
and
Assume first that for all By part (a) and by (2), the equivalent condition (2)(b) holds. It remains to show the fusion rules. But these follow from (2.1), (2.2) and (2.3).
Assume next that is a primitive axis of Jordan type . By (2), and by part (a), for all
Using Lemma 2.1 and the fusion rule we see that
Lemma 2.4.
Let be an axis. For write with and
Proof.
(1) Hence
(2) β
3. Derivations
We recall Definition 1.1. In particular, in this section is an axis.
Definition 3.1.
A derivation on is a -linear map that satisfies Leibnizβs law:
For we denote and so
Lemma 3.2.
Let and Write then
Proof.
(1)Β First note that hence
(2) By (1),
(3), (4), (5) and (6) follow from (2). β
Lemma 3.3.
Notation as in Lemma 3.2,
iff
Proof.
(1) This follows from Lemma 3.2(4), since and
(2) Using Lemma 3.2(4) we get,
(3) By Lemma 3.2(5),
(4) Follows from (1), (2) and (3). β
Lemma 3.4.
Notation as in Lemma 3.2,
iff
Proof.
(1) This follows from Lemma 3.2(3), since
(2) By Lemma 3.2(5),
(3) By Lemma 3.2(3),
(4) Follows from (1), (2) and (3).
β
Lemma 3.5.
Notation as in Lemma 3.2.
iff
Proof.
(1) By Lemma 3.2(4),
(2) By Lemma 3.2(2)
(3) By Lemma 3.2(6),
(4) Follows from (1), (2) and (3). β
Lemma 3.6.
Proof.
(1) By Lemma 3.2(3),
(2) By Lemma 3.2(4),
(3) By Lemma 3.2(6),
(4) This is immediate from (1), (2) and (3). β
Theorem 3.7.
Let be axes. Then the following are equivalent
is a derivation on
for all and
for all
4. Obtaining identities in
We recall Definition 1.1. In particular and are axes.
Lemma 4.1.
Let Then,
and
Proof.
(1) Write Then, by Lemma 2.4(2) applied to
(2) Replacing in (1) we get
where,
Since we see that
(3) Since is an idempotent we have
Since comparing the components of both sides yields (a) and (e) and Now by (2), so (c) holds. Since
(a) implies so (b) holds. Then, since we see that (d) holds.
In the proofs of (4) and (5) ahead, stands for the -component plus the -component of the right side of the equation.
(4) Replacing in (1) we get,
with and where,
where we used that and part (2) for and Since (4) follows.
(5) Replacing in (1) we get,
with where we used This shows part (b). Now
where the last equality follows from
We have
so part (a) holds. β
Lemma 4.2.
Let be axes in Then
for all that is
For all
Taking and in (2) yields
Taking and in (2), we get
Taking and in (2) we get
From (a) and (b) it follows:
Taking and in (2) we get
For all
Interchanging and in (7) we get
By subtracting we get
By adding we get
For all
Interchanging and in (9), we get
by subtracting,
By adding,
Proof.
(2) By (1),
where we used
(c) By Lemma 4.1(5)(b), and by (b),
(d) Note that Hence, by (c),
(c) We only need to compute (7) Comparing the -components in (6) we get
Lemma 4.3.
Let and then
Proof.
Lemma 4.4.
Let be an axis of Jordan type half. Then
and
and
Remark 4.5.
Lemma 4.6.
Putting in Lemma 4.4 gives
Proof.
(3) Using (2) and Lemma 4.1(4),
Theorem 4.7.
Let and then
Proof.
Lemma 4.8.
Putting and in Lemma 4.4 we get
Hence
Proof.
Hence
This shows (1) and similarly (4) holds.
Hence, by (2),
Multiplying by using and we get
Now we use to see that
so after cancellation we are left with β
Lemma 4.9.
Putting and in Lemma 4.4 we get
Proof.
By Lemma 4.4, replacing with and with
Hence, since the -component is
And, since the -component plus the -compoments are
Lemma 4.10.
For all
Proof.
5. Main theorems
In this section we prove our main results. As before, are axes.
Theorem 5.1.
If then
Proof.
We first show
We now complete the proof of the theorem. Interchanging in (5.1) and adding, we get
Theorem 5.2.
Assume then
For all
For all
If then for all
If then either or
Proof.
(1):Β Using Lemma 4.8(4) and Lemma 4.2(10b) we get
(using Theorem 5.1) | ||
(using Theorem 4.7(3)) | ||
(2): Follows from (1) and Lemma 4.6(3), since
(3): Let and Recall from Theorem 4.7(6) that and similarly for By (1) we get,
We conclude that
Multiplying (5.2) by and using (2) we see that
This is because by (2),
Assume Notice that since for all we see that with for Hence, by (5.3),
Assume furthermore that Taking (so that ) we see that since we conclude that
As a corollary, we get
Theorem 5.3.
Assume then
for all
Proof.
(1):Β By Theorem 4.7(4a), by Theorem 5.2 (3), Therefore, for some Multiplying by and recalling that we see that so and (1) hold.
(2): By Theorem 4.7 (8), Therefore, by (1),
(3):Β By Theorem 4.7(6), so, by (1),
(4):Β By Theorem 4.7(3), By (3), Therefore, β
Theorem 5.4.
If then
Proof.
Theorem 5.5.
If then
As an immediate corollary we have
Theorem 5.6.
If then is a derivation.
6. The case
In this section we assume that are axes and that We denote by the radical of the Frobenius form.
Theorem 6.1.
Proof.
First note that by [HSS, Lemma 3.1.9], By Lemma 4.2(7), Theorem 5.1, and Theorem 4.7(7), for all
Multiplying (6.1) by using Theorem 4.7 (1) and (8) we get
Replacing with in (6.1) using Theorem 4.7 (2) and (7) and (6.2), we get
Multiplying (6.3) by using (6.2) we now get
so the theorem holds. β
Theorem 6.2.
If then is a derivation.
Assume Then
Assume Then
In particular, is a derivation.
Proof.
Note also that
(2) (A) We first show that
Let By Lemma 4.10(1),
We now use Theorem 4.7 and By Theorem 4.7(1), and by Theorem 4.7(5), Thus, changing with in (6.6) we get,
By Lemma 4.10(2),
Putting in place of and using again that and (6.5), we get
So multiplying by yields Since (A) holds.
Lemma 6.3.
(6.9) -
Proof.
(1) Replacing in (6.6) by we get
Lemma 6.4.
We can now prove part (2B) of the theorem. By Lemma 4.9(2)
So part (Bi) follows from Lemma 6.4(4ii), since
(C) We start this part by showing
Lemma 6.5.
Proof.
Putting in place of in (6.7), we get
Now part (2C) follows from (6.7), since and as
7. The case
Recall that denotes the radical of the Frobenius form. In this section we prove
Theorem 7.1.
If then is a derivation.
We prove the theorem by showing that assertions (a), (b) and (c) of Theorem 3.7(ii) hold. Assertion (a) holds by Theorem 5.1. Theorems 7.2, 7.3 and 7.4 deal with assertions (b) and (c).
7.1. The case where
In this subsection we assume that Note that is an ideal. Note also, that if then because In particular, if then
Theorem 7.2.
Assume then
Proof.
By Lemma 4.10(2)
Replacing with in (7.1) using Theorem 4.7(5), we get
because by Theorem 4.7(1),
By Lemma 4.10(1),
Theorem 7.3.
Let and then
Proof.
Using also Theorem 4.7, we have
By Lemma 4.9(2)
and by Lemma 4.3(3)
We prove the theorem with a series of short claims.
CLAIM 1.
Proof.
CLAIM 2.
CLAIM 3.
Proof.
(1):Β Replacing in CLAIM 2, with we get
Hence
Multiplying by gives (1).
(2): Replacing in (7.10) with using and CLAIM 1(2),
CLAIM 4.
Proof.
(1):Β By CLAIM 3(1), and by Multiplying CLAIM 1(1) by we get
CLAIM 5.
Proof.
This follows from Lemma 4.9(1), Also using CLAIM 4(2). β
CLAIM 6.
Proof.
Multiplying (7.11) by using CLAIM 3(2),
CLAIM 7.
Proof.
This follows from CLAIM 2 and CLAIM 6. β
CLAIM 8.
Proof.
Part (1) follows from (7.10) and CLAIM 1(1), and (2) follows from (1) and CLAIM 6. β
Notice now that the theorem is immediate from CLAIM 7 and CLAIM 8(2). β
7.2. The case where
Theorem 7.4.
Let and then
Appendix A
In this appendix, we put two lemmas that are not being used in this paper but may find applications in the future.
Lemma A.1.
Let and
Proof.
Lemma A.2.
Let then
Comparing the -component in (1) we get
Interchanging and in (2) we get
By subtracting:
By adding:
Comparing the -components in (1) we get
Interchanging in (4) and subtracting we get,
Interchanging in (4) and adding we get