### v6

# Two axes in non-associative algebras with a Frobenius form

Abstract.
Throughout this paper $A$ is a commutative non-associative algebra over a field $\mathbb{F}$ of characteristic not $2.$ In addition $A$ possesses a Frobenius form. We obtain detailed information about the multiplication in $A$ given two axes of type half in $A.$
Key words and phrases:
idempotent, axis, primitive, Jordan type, axial algebra, derivation
2020 Mathematics Subject Classification:
Primary: 17C27, 17A99, Secondary: 17C10
1. Introduction
Throughout this paper $A$ is a commutative non-associative algebra over a field $\mathbb{F}$ of characteristic not $2.$ Assume that (,)(\ ,\ ) is a Frobenius form on $A,$ that is, (,)(\ ,\ ) is a symmetric bilinear map such that $(xy,z)=(x,yz),$ $\forall x,y,z\in A.$ We denote by $\mathfrak{R}$ the radical of (,)(\ ,\ ). Recall that for $z\in A,\ \lambda\in\mathbb{F},$ we denote $A_{\lambda}(z)=\{x\in A\mid zx=\lambda x\}.$ We direct the reader to Definition 1.1 below and the beginning of §2 for the definitions required for the introduction.
This paper was inspired by [D], which in turn was inspired by [GSS]. These two papers reveal an unexpected, and rather amazing structure in primitive axial algebras of Jordan type half. Since axial algebras are related to $3$-transposition groups, on the one hand, and to the Griess Monster algebra, on the other hand, there must be something real there (:)).
Now when one looks carefully on the arguments in [D], one realizes that all that is being used for an idempotent $c\in A$ is that: $P_{c}(x,y)=0,\ \forall x,y\in A$ iff $c$ is an axis (where in characteristic $3$ one has to also require $(c,c)=1,$ see Lemma 2.3). However, some of the arguments in [D] are sleek, smart (and short), but I thought there should be more elementary (albeit longer) ones. Thus, I decided to embark on the tedious (though rewarding) effort to make the arguments elementary, and, on the way, there are some very interesting identities that are obtained.
Indeed, Theorem 1.4 introduces new identities, and there are many additional identities in this paper which we believe will become very useful, in particular, for giving an elementary proof that finitely generated three transposition groups are finite. Theorem 1.2 is new as well.
Definitions and notation 1.1.
  1. (1)
    Recall that an idempotent $c\in A$ is called a primitive axis of Jordan type $\frac{1}{2},$ if $A=\mathbb{F}c\oplus A_{0}(c)\oplus A_{\frac{1}{2}}(c),$ and $(c,c)=1.$ Further, $c$ satisfies the following fusion rules: for $x_{0},y_{0}\in A_{0}(c)$ and $x_{\frac{1}{2}},y_{\frac{1}{2}}\in A_{\frac{1}{2}}(c),$ $x_{0}y_{0}\in A_{0}(c),\ x_{0}y_{\frac{1}{2}}\in A_{\frac{1}{2}}(c),$ and $x_{\frac{1}{2}}y_{\frac{1}{2}}\in\mathbb{F}c+A_{0}(c).$ For short we will say that $c$ is an axis.
  2. (2)
    Throughout the paper, $a,b$ are fixed axes in $A.$
  3. (3)
    Throughout the paper, except for §2, when we write $x=\alpha_{x}a+x_{0}+x_{\frac{1}{2}},\ x\in A,$ we mean $\alpha_{x}\in\mathbb{F},\ x_{\lambda}\in A_{\lambda}(a),\ \lambda\in\{0,\frac{1}{2}\}.$ Lemma 2.1 below shows that $\alpha_{x}=(a,x).$
  4. (4)
    Throughout the paper, except for §2, when we write $x_{\lambda},\ \lambda\in\{0,\frac{1}{2}\},$ (but there is no element $x$ around) we just mean $x_{\lambda}\in A_{\lambda}(a).$
  5. (5)
    Recall that for an axis $a$ the Miyamoto map $\tau_{a}:x\mapsto(a,x)a+x_{0}-x_{\frac{1}{2}},$ where $x=(a,x)a+x_{0}+x_{\frac{1}{2}}\in A,$ is an automorphism of order at most $2$ of $A.$
Theorem 1.2 (Theorem 3.7).
Let $a,b$ be axes. Then the following are equivalent
  1. (i)
    $L_{a},L_{b}]$ is a derivation on $A.$
  2. (ii)
    1. (a)
      $b_{\frac{1}{2}}(x_{0}y_{0})=(b_{\frac{1}{2}}x_{0})y_{0}+(b_{\frac{1}{2}}y_{0})x_{0},\ \forall x_{0},y_{0}\in A_{0}(a).$
    2. (b)
      $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}=(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0},$ for all $x_{\frac{1}{2}}\in A_{\frac{1}{2}}(a)$ and $y_{0}\in A_{0}(a).$
    3. (c)
      $b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}=$
      $\frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}},$
      for all $x_{\frac{1}{2}},y_{\frac{1}{2}}\in A_{\frac{1}{2}}(a).$
The following result is proven in §5, and is essentially a consequence of [D, Theorem 1.2 and Proposition 4.3], but here we prove it using elementary means.
Theorem 1.3 (Theorem 5.1).
If $(a,b)\neq\frac{1}{4},$ then $b_{\frac{1}{2}}(x_{0}y_{0})=(b_{\frac{1}{2}}x_{0})y_{0}+(b_{\frac{1}{2}}y_{0})x_{0}.$
Theorem 1.4 (Theorem 4.9).
Let $x_{0}\in A_{0}(a)$ and $x_{\frac{1}{2}}\in A_{\frac{1}{2}}(a),$ then
  1. (1)
    $(x_{\frac{1}{2}}b_{0})b_{0}=\frac{1}{2}(1-(a,b))x_{\frac{1}{2}}b_{0}.$
  2. (2)
    $(x_{0}b_{0})b_{0}=\frac{1}{2}(1-(a,b))x_{0}b_{0}+\frac{1}{2}(b,x_{0})b_{0}.$
  3. (3)
    $(x_{0}b_{\frac{1}{2}})b_{\frac{1}{2}}=\frac{1}{2}(a,b)(b,x_{0})a+\frac{1}{2}(a,b)x_{0}b_{0}.$
  4. (4)
    1. (a)
      $\Big(x_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}b_{\frac{1}{2}}=\frac{1}{4}(a,b)(1-(a,b))x_{\frac{1}{2}}-\frac{1}{2}(a,b)x_{\frac{1}{2}}b_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}.$
    2. (b)
      $(x_{\frac{1}{2}}b_{\frac{1}{2}})b_{\frac{1}{2}}=\frac{1}{4}(a,b)(1-(a,b))x_{\frac{1}{2}}-\frac{1}{2}(a,b)x_{\frac{1}{2}}b_{0}+\frac{1}{2}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}.$
  5. (5)
    $(x_{\frac{1}{2}}b_{0})b_{\frac{1}{2}}=\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}})a+\frac{1}{4}(b,x_{\frac{1}{2}})b_{0}.$
  6. (6)
    $(x_{\frac{1}{2}}b_{\frac{1}{2}})b_{0}=\frac{1}{2}(1-(a,b))\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})b_{0}.$
  7. (7)
    $(x_{0}b_{0})b_{\frac{1}{2}}=\frac{1}{2}(1-(a,b))b_{\frac{1}{2}}x_{0}+\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}.$
  8. (8)
    $(x_{0}b_{\frac{1}{2}})b_{0}=\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}.\quad$
Theorem 1.5 (which infact are two separtate theorems) is proven in §6 and §7 by elemntary means, compare with [D, Theorem 1.2 and Proposition 4.3]).
Theorem 1.5 (Theorems 6.1 and 7.1 respectively).
Suppose $(a,b)\notin\{1,\frac{1}{4}\},$ then
  1. (1)
    $b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}=$
    $\frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
  2. (2)
    $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}=(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}.$
In particular, $L_{a},L_{b}]$ is a derivation.
For case $(a,b)=1$, we have the following theorem, proven in §8 by elemntary means.
Theorem 1.6.
Assume $(a,b)=1.$ Then
  1. (1)
    If $b_{0}=0,$ then $L_{a},L_{b}]$ is a derivation.
  2. (2)
    If ${\rm char}(\mathbb{F})\neq 3,$ then idetities (1) and (2) of Theorem 1.5 hold. In particular, $L_{a},L_{b}]$ is a derivation.
We left the case $(a,b)=\frac{1}{4}$ for future works.
2. The polynomials $Q_{c}(x)$ and $P_{c}(x,y)$
Recall Definition 1.1(2). Let $c\in A$ be an idempotent. Define, as in [D],
${Q_{c}(x):=c(cx)-\frac{1}{2}(cx+(c,x)c),}$
$P_{c}(x,y):=4(cx)(cy)-(c,y)cx-(cy)x-(c,x)cy-(cx)y-(c,xy)c+c(xy).$
Lemma 2.1.
Let $z\in A,\ \lambda,\delta\in\mathbb{F},\ x_{\lambda}\in A_{\lambda}(z),\ y_{\delta}\in A_{\delta}(z).$ If $\delta\neq\lambda,$ then $(x_{\lambda},y_{\delta})=0.$ In particular $x=(a,x)a+x_{0}+x_{\frac{1}{2}},$ where $x\in A,\ x_{0}\in A_{0}(a),\ x_{\frac{1}{2}}\in A_{\frac{1}{2}}(a).$
Proof.
If $\lambda=0,$ then, by hypothesis, $\delta\neq 0,$ and $(x_{\lambda},y_{\delta})=(x_{\lambda},\frac{1}{\delta}zy_{\delta})=\frac{1}{\delta}(zx_{\lambda},y_{\delta})=\frac{1}{\delta}(0,y_{\delta})=0.$ Assume $0\notin\{\lambda,\delta\}.$ Then
${(x_{\lambda},y_{\delta})=(\frac{1}{\lambda}zx_{\lambda},\ y_{\delta})=(\frac{1}{\lambda}x_{\lambda},\ zy_{\delta})=\frac{\delta}{\lambda}(x_{\lambda},y_{\delta}).}$
Hence if $\lambda\neq\delta,$ then $(x_{\lambda},y_{\delta})=0.$
Lemma 2.2 (Serss’ Lemma).
Let $a$ be an axis. Then $a(xy)=(ax)y,$ for all $x\in A$ and $y\in\mathbb{F}a+A_{0}(a).$
Proof.
Write $y=(a,y)a+y_{0},$ and $x=(a,x)a+x_{0}+x_{\frac{1}{2}},$ with $x_{0},y_{0}\in A_{0}(a)$ and $x_{\frac{1}{2}}\in A_{\frac{1}{2}}(a).$ Then
${xy=(a,x)(a,y)a+x_{0}y_{0}+\frac{1}{2}(a,y)x_{\frac{1}{2}}+x_{\frac{1}{2}}y_{0},}$
so
${a(xy)=(a,x)(a,y)a+\frac{1}{4}(a,y)x_{\frac{1}{2}}+\frac{1}{2}x_{\frac{1}{2}}y_{0},}$
while
${(ax)y=((a,x)a+\frac{1}{2}x_{\frac{1}{2}})((a,y)a+y_{0})=(a,x)(a,y)a+\frac{1}{4}(a,y)x_{\frac{1}{2}}+\frac{1}{2}x_{\frac{1}{2}}y_{0}.}\qed$
Lemma 2.3.
Let $c\in A$ be an idempotent, then
  1. (1)
    If ${\rm char}(\mathbb{F})\neq 3,$ and $P_{c}(c,c)=0,$ then $(c,c)=1.$
  2. (2)
    The following conditions on $c$ are equivalent:
    1. (a)
      $Q_{c}(x)=0,$ for all $x\in A.$
    2. (b)
      $A=\mathbb{F}c\oplus A_{0}(c)\oplus A_{\frac{1}{2}}(c),$ and $(c,c)=1.$
  3. (3)
    Assume $(c,c)=1.$ Then
    1. (a)
      $P_{c}(c,y)=4Q_{c}(y).$
    2. (b)
      $c$ is a primitive axis of Jordan type $\frac{1}{2}$ iff $P_{c}(x,y)=0,$ for all $x,y\in A.$
Proof.
(1): By definition, $P_{c}(c,c)=4c-(c,c)c-c-(c,c)c-c-(c,c)c+c=3c-3(c,c)c,$ so (1) holds.
(2) $(a)\implies(b):$ Assume that $Q_{c}(x)=0,$ for all $x\in A.$ Then $cx-\frac{1}{2}x-\frac{1}{2}(c,x)c\in A_{0}(c)$, and c(2cx2(c,x)c)=2(c(cx))2(c,x)c)=cx+(c,x)c2(c,x)c=cx(c,x)c.c(2cx-2(c,x)c)=2(c(cx))-2(c,x)c)=cx+(c,x)c-2(c,x)c=cx-(c,x)c. Thus $2cx-2(c,x)c\in A_{\frac{1}{2}}(c).$ Also $x=(c,x)c-2(cx-\frac{1}{2}x-\frac{1}{2}(c,x)c)+2cx-2(c,x)c.$ Hence, $A=\mathbb{F}c\oplus A_{0}(c)\oplus A_{\frac{1}{2}}(c).$
Since $0=Q_{c}(c)=\frac{1}{2}c-\frac{1}{2}(c,c)c,$ we see that $(c,c)=1.$
$(b)\implies(a):$ Assume that $(b)$ holds. Since $(c,c)=1,$ $Q_{c}(c)=0.$ Let $x_{0}\in A_{0}(c)$ and $x_{\frac{1}{2}}\in A_{\frac{1}{2}}(c).$ By Lemma 2.1, $(c,x_{0})=(c,x_{\frac{1}{2}})=0.$ Now $Q_{c}(x_{0})=c(cx_{0})-\frac{1}{2}(cx_{0})-\frac{1}{2}(c,x_{0})c)=0,$ and $Q_{c}(x_{\frac{1}{2}})=c(cx_{\frac{1}{2}})-\frac{1}{2}cx_{\frac{1}{2}}+(c,x_{\frac{1}{2}})c)=\frac{1}{4}x_{\frac{1}{2}}-\frac{1}{4}x_{\frac{1}{2}}-\frac{1}{2}(c,x_{\frac{1}{2}})c=0.$
(3)(a): We have
$P_{c}(c,y)$ $=4c(cy)-(c,y)c-(cy)c-(c,c)cy-cy-(c,cy)c+c(cy)$
$=4c(cy)-2cy-2(c,y)c=4Q_{c}(y),$
because $(c,c)=1,$ and $(c,cy)=(c,y),$ since (,)(\ ,\ ) is Frobenius.
(b): Assume that $A=\mathbb{F}c\oplus A_{0}(c)\oplus A_{\frac{1}{2}}(c).$ Let $x_{0},y_{0}\in A_{0}(c)$ and $x_{\frac{1}{2}},y_{\frac{1}{2}}\in A_{\frac{1}{2}}(c).$ By Lemma 2.1, $(c,x_{0})=(c,y_{0})=(c,x_{\frac{1}{2}})=(c,y_{\frac{1}{2}})=0.$ Also since (,)(\ ,\ ) is Frobenius, $(c,x_{0}y_{0})=0=(c,x_{0}y_{\frac{1}{2}}).$ Hence
(2.1) $P_{c}(x_{0},y_{0})$ $=4(cx_{0})(cy_{0})-(c,y_{0})cx_{0}-(cy_{0})x_{0}$
$-(c,x_{0})cy_{0}-(cx_{0})y_{0}-(c,x_{0}y_{0})c+c(x_{0}y_{0})$
$=0-0-0-0-0-0+c(x_{0}y_{0})$
$=c(x_{0}y_{0}),$
(2.2) $P_{c}(x_{0},y_{\frac{1}{2}})$ $=4(cx_{0})(cy_{\frac{1}{2}})-(c,y_{\frac{1}{2}})cx_{0}-(cy_{\frac{1}{2}})x_{0}-(c,x_{0})cy_{\frac{1}{2}}$
$-(cx_{0})y_{\frac{1}{2}}-(c,x_{0}y_{\frac{1}{2}})c+c(x_{0}y_{\frac{1}{2}})$
$=\textstyle{0-0-\frac{1}{2}x_{0}y_{\frac{1}{2}}-0-0-0+c(x_{0}y_{\frac{1}{2}})}$
$=\textstyle{-\frac{1}{2}x_{0}y_{\frac{1}{2}}+c(x_{0}y_{\frac{1}{2}}),}$
and
(2.3) $P_{c}(x_{\frac{1}{2}},y_{\frac{1}{2}})$ $=4(cx_{\frac{1}{2}})(cy_{\frac{1}{2}})-(c,y_{\frac{1}{2}})cx_{\frac{1}{2}}-(cy_{\frac{1}{2}})x_{\frac{1}{2}}$
$-(c,x_{\frac{1}{2}})cy_{\frac{1}{2}}-(cx_{\frac{1}{2}})y_{\frac{1}{2}}-(c,x_{\frac{1}{2}}y_{\frac{1}{2}})c+c(x_{\frac{1}{2}}y_{\frac{1}{2}})$
$=\textstyle{x_{\frac{1}{2}}y_{\frac{1}{2}}-0-\frac{1}{2}y_{\frac{1}{2}}x_{\frac{1}{2}}-0-\frac{1}{2}x_{\frac{1}{2}}y_{\frac{1}{2}}}$
$-(c,x_{\frac{1}{2}}y_{\frac{1}{2}})c+c(x_{\frac{1}{2}}y_{\frac{1}{2}})$
$=-(c,x_{\frac{1}{2}}y_{\frac{1}{2}})c+c(x_{\frac{1}{2}}y_{\frac{1}{2}}).$
Assume first that $P_{c}(x,y)=0,$ for all $x,y\in A.$ By part (a) and by (2), the equivalent condition (2)(b) holds. It remains to show the fusion rules. But these follow from (2.1), (2.2) and (2.3).
Assume next that $c$ is a primitive axis of Jordan type $\frac{1}{2}$. By (2), and by part (a), $P_{c}(c,y)=P_{c}(x,c)=0,$ for all $x,y\in A.$
Using Lemma 2.1 and the fusion rule $x_{\frac{1}{2}}y_{\frac{1}{2}}\in\mathbb{F}c+A_{0}(c),$ we see that $c(x_{\frac{1}{2}}y_{\frac{1}{2}})=(c,x_{\frac{1}{2}}y_{\frac{1}{2}})c.$
Now, by the fusion rules and by (2.1), (2.2) and (2.3), $P(x_{0},y_{0})=P(x_{0},y_{\frac{1}{2}})=P(y_{\frac{1}{2}},x_{0})=P(x_{\frac{1}{2}},y_{\frac{1}{2}})=0,$ for all $x_{0},y_{0}\in A_{0}(c)$ and $x_{\frac{1}{2}},y_{\frac{1}{2}}\in A_{\frac{1}{2}}(c).$ Hence, $P(x,y)=0,$ for all $x,y\in A.$
The following lemma will be used in the next section.
Lemma 2.4.
Let $a\in A$ be an axis. For $x\in A,$ write $x=(a,x)a+x_{0}+x_{\frac{1}{2}},$ with $x_{0}\in A_{0}(a)$ and $x_{\frac{1}{2}}\in A_{\frac{1}{2}}(a).$
  1. (1)
    $x_{0}=x-2ax+(a,x)a.$
  2. (2)
    $x_{\frac{1}{2}}=2ax-2(a,x)a.$
Proof.
(1) $ax=(a,x)a+\frac{1}{2}x_{\frac{1}{2}}.$ Hence $x+(a,x)a-2ax=(a,x)a+x_{0}+x_{\frac{1}{2}}+(a,x)a-2(a,x)a-x_{\frac{1}{2}}=x_{0}.$
(2) $2ax-2(a,x)a=2(a,x)a+x_{\frac{1}{2}}-2(a,x)a=x_{\frac{1}{2}}.$
3. Derivations
We recall Definition 1.1. In particular, in this section $a$ is an axis.
Definition 3.1.
  1. (1)
    A derivation on $A$ is a $\mathbb{F}$-linear map $D:A\to A$ that satisfies Leibniz’s law: $D(xy)=xD(y)+D(x)y.$
  2. (2)
    For $x,y\in A$ we denote $L_{x}(y)=xy,$ and $L_{x},L_{y}]=L_{x}L_{y}-L_{y}L_{x},$ so $L_{x},L_{y}](z)=x(yz)-y(xz).$
Lemma 3.2.
Let $v\in A_{\frac{1}{2}}(a)$ and $x\in A.$ Write $x=(a,x)a+x_{0}+x_{\frac{1}{2}},$ then
  1. (1)
    $L_{a},L_{v}](x)=-2v(ax)+\frac{1}{2}(vx+(v,x)a+(a,x)v).$
  2. (2)
    $L_{a},L_{v}](x)=\frac{1}{2}(v,x)a-\frac{1}{2}vx_{\frac{1}{2}}-\frac{1}{4}(a,x)v+\frac{1}{2}vx_{0}.$
  3. (3)
    $\Big([L_{a},L_{v}](x)\Big)_{0}=-\frac{1}{2}\Big(vx_{\frac{1}{2}}\Big)_{0}.$
  4. (4)
    $\Big([L_{a},L_{v}](x)\Big)_{\frac{1}{2}}=\frac{1}{2}vx_{0}-\frac{1}{4}(a,x)v.$
  5. (5)
    $L_{a},L_{v}](x_{0})=\frac{1}{2}(vx_{0}).$
  6. (6)
    $L_{a},L_{v}](x_{\frac{1}{2}})=\frac{1}{2}(v,x_{\frac{1}{2}})a-\frac{1}{2}vx_{\frac{1}{2}}.$
Proof.
(1) First note that $ax=(a,x)a+\frac{1}{2}x_{\frac{1}{2}}=\frac{1}{4}(x-x^{\tau_{a}})+(a,x)a,$ hence
$a(vx)$ $=\textstyle{\frac{1}{4}(vx-(vx)^{\tau_{a}})+(a,vx)a=\frac{1}{4}(vx+vx^{\tau_{a}})+\frac{1}{2}(v,x)a}$
$=\textstyle{\frac{1}{4}(vx+v(x+4(a,x)a-4(ax)))+\frac{1}{2}(v,x)a}$
$=\textstyle{\frac{1}{4}(vx+vx+2(a,x)v-4v(ax))+\frac{1}{2}(v,x)a}$
$=\textstyle{-v(ax)+\frac{1}{2}(vx+(v,x)a+(a,x)v).}$
(2) By (1),
$[L_{a},L_{v}](x)$ $=\textstyle{-2v(ax)+\frac{1}{2}(vx+(v,x)a+(a,x)v)}$
$=\textstyle{-2v((a,x)a+\frac{1}{2}x_{\frac{1}{2}})+\frac{1}{2}(\frac{1}{2}(a,x)v+vx_{0}+vx_{\frac{1}{2}}+(v,x)a+(a,x)v)}$
$=\textstyle{-(a,x)v-vx_{\frac{1}{2}}+\frac{3}{4}(a,x)v+\frac{1}{2}vx_{0}+\frac{1}{2}vx_{\frac{1}{2}}+\frac{1}{2}(v,x)a}$
$=\textstyle{\frac{1}{2}(v,x)a-\frac{1}{2}vx_{\frac{1}{2}}-\frac{1}{4}(a,x)v+\frac{1}{2}vx_{0}.}$
(3), (4), (5) and (6) follow from (2). ∎
Lemma 3.3.
Notation as in Lemma 3.2,
  1. (1)
    $\Big([L_{a},L_{v}](xy_{0})\Big)_{\frac{1}{2}}=\frac{1}{2}v(y_{0}x_{0}).$
  2. (2)
    $\Big(\big([L_{a},L_{v}](x)\big)y_{0}\Big)_{\frac{1}{2}}=\frac{1}{2}(vx_{0})y_{0}-\frac{1}{4}(a,x)vy_{0}.$
  3. (3)
    $\Big(\big([L_{a},L_{v}](y_{0})\big)x\Big)_{\frac{1}{2}}=\frac{1}{2}(vy_{0})x_{0}+\frac{1}{4}(a,x)vy_{0}.$
  4. (4)
    $\Big([L_{a},L_{v}](xy_{0})\Big)_{\frac{1}{2}}=\Big(([L_{a},L_{v}](x))y_{0}\Big)_{\frac{1}{2}}+\Big(([L_{a},L_{v}](y_{0}))x\Big)_{\frac{1}{2}},$ iff
    $v(x_{0}y_{0})=(vx_{0})y_{0}+(vy_{0})x_{0}.$
Proof.
(1) This follows from Lemma 3.2(4), since $(a,xy_{0})=0,$ and $\Big(xy_{0}\Big)_{0}=x_{0}y_{0}.$
(2) Using Lemma 3.2(4) we get,
${\Big(\big([L_{a},L_{v}](x)\big)y_{0}\Big)_{\frac{1}{2}}=\Big([L_{a},L_{v}](x)\Big)_{\frac{1}{2}}y_{0}=\frac{1}{2}(vx_{0})y_{0}-\frac{1}{4}(a,x)vy_{0}.}$
(3) By Lemma 3.2(5), $\Big(\big([L_{a},L_{v}](y_{0})\big)x\Big)_{\frac{1}{2}}=\frac{1}{2}\Big((vy_{0})x\Big)_{\frac{1}{2}}=\\ \frac{1}{2}((vy_{0})x_{0}+\frac{1}{2}(a,x)vy_{0})=\frac{1}{2}(vy_{0})x_{0}+\frac{1}{4}(a,x)vy_{0}.$
(4) Follows from (1), (2) and (3). ∎
Lemma 3.4.
Notation as in Lemma 3.2,
  1. (1)
    $\Big([L_{a},L_{v}](xy_{0})\Big)_{0}=-\frac{1}{2}\Big(v(x_{\frac{1}{2}}y_{0})\Big)_{0}.$
  2. (2)
    $\Big(\big([L_{a},L_{v}](y_{0})\big)x\Big)_{0}=\frac{1}{2}\Big((vy_{0})x_{\frac{1}{2}}\Big)_{0}.$
  3. (3)
    $\Big(\big([L_{a},L_{v}](x)\big)y_{0}\Big)_{0}=-\frac{1}{2}(vx_{\frac{1}{2}})y_{0}.$
  4. (4)
    $\Big([L_{a},L_{v}](xy_{0})\Big)_{0}=\Big([L_{a},L_{v}](x)\Big)_{0}y_{0}+\Big([L_{a},L_{v}](y_{0})x\Big)_{0},$ iff
    $\Big(v(x_{\frac{1}{2}}y_{0})\Big)_{0}=(vx_{\frac{1}{2}})y_{0}-\Big((vy_{0})x_{\frac{1}{2}}\Big)_{0}.$
Proof.
(1) This follows from Lemma 3.2(3), since $\Big(xy_{0}\Big)_{\frac{1}{2}}=x_{\frac{1}{2}}y_{0}.$
(2) By Lemma 3.2(5), $\Big(\big([L_{a},L_{v}](y_{0})\big)x\Big)_{0}=\frac{1}{2}\Big((vy_{0})x\Big)_{0}=\frac{1}{2}\Big((vy_{0})x_{\frac{1}{2}}\Big)_{0}.$
(3) By Lemma 3.2(3), $\Big(\big([L_{a},L_{v}](x)\big)y_{0}\Big)_{0}=\Big([L_{a},L_{v}](x)\Big)_{0}y_{0}=-\frac{1}{2}\Big(vx_{\frac{1}{2}}\Big)_{0}y_{0}=-\frac{1}{2}(vx_{\frac{1}{2}})y_{0}.$
(4) Follows from (1), (2) and (3).
Lemma 3.5.
Notation as in Lemma 3.2.
  1. (1)
    $\Big([L_{a},L_{v}](xy_{\frac{1}{2}})\Big)_{\frac{1}{2}}=\frac{1}{2}v(x_{\frac{1}{2}}y_{\frac{1}{2}})-\frac{1}{2}(a,x_{\frac{1}{2}}y_{\frac{1}{2}})v.$
  2. (2)
    $\Big(\big([L_{a},L_{v}](x)\big)y_{\frac{1}{2}}\Big)_{\frac{1}{2}}=\frac{1}{4}(v,x)y_{\frac{1}{2}}-\frac{1}{2}(vx_{\frac{1}{2}})y_{\frac{1}{2}}.$
  3. (3)
    $\Big(\big([L_{a},L_{v}](y_{\frac{1}{2}})\big)x\Big)_{\frac{1}{2}}=\frac{1}{4}(v,y)x_{\frac{1}{2}}-\frac{1}{2}(vy_{\frac{1}{2}})x_{\frac{1}{2}}.$
  4. (4)
    $\Big([L_{a},L_{v}](xy_{\frac{1}{2}})\Big)_{\frac{1}{2}}=\Big(\big([L_{a},L_{v}](x)\big)y_{\frac{1}{2}}\Big)_{\frac{1}{2}}+\Big(\big([L_{a},L_{v}](y_{\frac{1}{2}})\big)x\Big)_{\frac{1}{2}},$ iff
    ${v(x_{\frac{1}{2}}y_{\frac{1}{2}})+(vx_{\frac{1}{2}})y_{\frac{1}{2}}+(vy_{\frac{1}{2}})x_{\frac{1}{2}}=\frac{1}{2}(v,x)y_{\frac{1}{2}}+\frac{1}{2}(v,y)x_{\frac{1}{2}}+(a,x_{\frac{1}{2}}y_{\frac{1}{2}})v.}$
Proof.
(1) By Lemma 3.2(4),
$\textstyle{\Big([L_{a},L_{v}](xy_{\frac{1}{2}})\Big)_{\frac{1}{2}}=\frac{1}{2}v\Big(xy_{\frac{1}{2}}\Big)_{0}-\frac{1}{4}(a,xy_{\frac{1}{2}})v=\frac{1}{2}v\Big(x_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}-\frac{1}{4}(a,x_{\frac{1}{2}}y_{\frac{1}{2}})v}$
$=\textstyle{\frac{1}{2}v(x_{\frac{1}{2}}y_{\frac{1}{2}})-\frac{1}{2}(a,x_{\frac{1}{2}}y_{\frac{1}{2}})av-\frac{1}{4}(a,x_{\frac{1}{2}}y_{\frac{1}{2}})v=\frac{1}{2}v(x_{\frac{1}{2}}y_{\frac{1}{2}})-\frac{1}{2}(a,x_{\frac{1}{2}}y_{\frac{1}{2}})v.}$
(2) By Lemma 3.2(2),
$\Big(\big([L_{a},L_{v}](x)\big)y_{\frac{1}{2}}\Big)_{\frac{1}{2}}=(\frac{1}{2}(v,x)a-\frac{1}{2}vx_{\frac{1}{2}})y_{\frac{1}{2}}=\frac{1}{4}(v,x)y_{\frac{1}{2}}-\frac{1}{2}(vx_{\frac{1}{2}})y_{\frac{1}{2}}.$
(3) By Lemma 3.2(6),
${\Big([L_{a},L_{v}](y_{\frac{1}{2}})x\Big)_{\frac{1}{2}}=(\frac{1}{2}(v,y_{\frac{1}{2}})a-\frac{1}{2}vy_{\frac{1}{2}})x_{\frac{1}{2}}=\frac{1}{4}(v,y)x_{\frac{1}{2}}-\frac{1}{2}(vy_{\frac{1}{2}})x_{\frac{1}{2}}.}$
(4) Follows from (1), (2) and (3). ∎
Lemma 3.6.
Notation as in Lemma 3.2,
  1. (1)
    $\Big([L_{a},L_{v}](xy_{\frac{1}{2}})\Big)_{0}=-\frac{1}{4}(a,x)\Big(vy_{\frac{1}{2}}\Big)_{0}-\frac{1}{2}\Big(v(x_{0}y_{\frac{1}{2}})\Big)_{0}$
  2. (2)
    $\Big(\big([L_{a},L_{v}](x)\big)y_{\frac{1}{2}}\Big)_{0}=\frac{1}{2}\Big((vx_{0})y_{\frac{1}{2}}\Big)_{0}-\frac{1}{4}(a,x)\Big(vy_{\frac{1}{2}}\Big)_{0}.$
  3. (3)
    $\Big(\big([L_{a},L_{v}](y_{\frac{1}{2}})\big)x\Big)_{0}=-\frac{1}{2}(vy_{\frac{1}{2}})x_{0}.$
  4. (4)
    $\Big([L_{a},L_{v}](xy_{\frac{1}{2}})\Big)_{0}=\Big(\big([L_{a},L_{v}](x)\big)y_{\frac{1}{2}}\Big)_{0}+\Big(\big([L_{a},L_{v}](y_{\frac{1}{2}})\big)x\Big)_{0},$ iff
    $\Big(v(y_{\frac{1}{2}}x_{0})\Big)_{0}=(vy_{\frac{1}{2}})x_{0}-\Big((vx_{0})y_{\frac{1}{2}}\Big)_{0}.$
    (Note that is exactly the same condition of Lemma 3.4(4).)
Proof.
(1) By Lemma 3.2(3),
$\textstyle{\Big([L_{a},L_{v}](xy_{\frac{1}{2}})\Big)_{0}=-\frac{1}{2}\Big(v\Big(xy_{\frac{1}{2}}\Big)_{\frac{1}{2}}\Big)_{0}=}$
$\textstyle{-\frac{1}{2}\Big(v((a,x)a+x_{0})y_{\frac{1}{2}})\Big)_{0}=-\frac{1}{4}(a,x)\Big(vy_{\frac{1}{2}}\Big)_{0}-\frac{1}{2}\Big(v(x_{0}y_{\frac{1}{2}})\Big)_{0}.}$
(2) By Lemma 3.2(4),
$\Big(\big([L_{a},L_{v}](x)\big)y_{\frac{1}{2}}\Big)_{0}=\Big(\Big([L_{a},L_{v}](x)\big)\Big)_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}$
$=\textstyle{\Big((\frac{1}{2}vx_{0}-\frac{1}{4}(a,x)v)y_{\frac{1}{2}}\Big)_{0}=\frac{1}{2}\Big((vx_{0})y_{\frac{1}{2}}\Big)_{0}-\frac{1}{4}(a,x)\Big(vy_{\frac{1}{2}}\Big)_{0}.}$
(3) By Lemma 3.2(6),
$\textstyle{\Big(\big([L_{a},L_{v}](y_{\frac{1}{2}})\big)x\Big)_{0}=\Big((\frac{1}{2}(v,y_{\frac{1}{2}})a-\frac{1}{2}vy_{\frac{1}{2}})x\Big)_{0}}$
$=\textstyle{-\frac{1}{2}\Big((vy_{\frac{1}{2}})x\Big)_{0}=-\frac{1}{2}(vy_{\frac{1}{2}})x_{0}.}$
(4) This is immediate from (1), (2) and (3). ∎
Theorem 3.7.
Let $a,b$ be axes. Then the following are equivalent
  1. (i)
    $L_{a},L_{b}]$ is a derivation on $A.$
  2. (ii)
    1. (a)
      $b_{\frac{1}{2}}(x_{0}y_{0})=(b_{\frac{1}{2}}x_{0})y_{0}+(b_{\frac{1}{2}}y_{0})x_{0},\ \forall x_{0},y_{0}\in A_{0}(a).$
    2. (b)
      $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}=(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0},$ for all $x_{\frac{1}{2}}\in A_{\frac{1}{2}}(a)$ and $y_{0}\in A_{0}(a).$
    3. (c)
      $b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}=$
      $\frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}},$
      for all $x_{\frac{1}{2}},y_{\frac{1}{2}}\in A_{\frac{1}{2}}(a).$
Proof.
First, note that
(3.1) $L_{a},L_{x_{0}}]=0,\text{ for all }x_{0}\in A_{0}(a),$
because by Seress’ Lemma 2.2, $L_{a},L_{x_{0}}](x)=a(x_{0}x)-x_{0}(ax)=0,$ for all $x\in A.$ Hence, since $L_{a},L_{a}]=0,$ we see that for $x\in A,$ $L_{a},L_{x}]$ is a derivation iff $L_{a},L_{x_{\frac{1}{2}}}]$ is a derivation. Let $v\in A_{\frac{1}{2}}(a).$ By Lemma 3.2(5&6),
$\textstyle{[L_{a},L_{v}](a)=a(av)-v(aa)=-\frac{1}{4}v,}$
$\textstyle{[L_{a},L_{v}](x_{\frac{1}{2}})=a(vx_{\frac{1}{2}})-\frac{1}{2}vx_{\frac{1}{2}},\text{ and }[L_{a},L_{v}](x_{0})=\frac{1}{2}vx_{0}.}$
Hence,
$\textstyle{\big([L_{a},L_{v}](x_{0})\big)a+\big([L_{a},L_{v}](a)\big)x_{0}=\frac{1}{4}vx_{0}-\frac{1}{4}vx_{0}=0=[L_{a},L_{v}](ax_{0}).}$
Also,
$\textstyle{\big([L_{a},L_{v}](a)\big)x_{\frac{1}{2}}+a\big([L_{a},L_{v}](x_{\frac{1}{2}})\big)=-\frac{1}{4}vx_{\frac{1}{2}}+\frac{1}{2}a(vx_{\frac{1}{2}})=[L_{a},L_{v}](ax_{\frac{1}{2}}).}$
Finally,
${[L_{a},L_{v}](a)=-\frac{1}{8}v-\frac{1}{8}v=a\big([L_{a},L_{v}](a)\big)+a\big([L_{a},L_{v}](a)\big)=[L_{a},L_{v}](aa).}$
Thus $L_{a},L_{v}](xa)=x[L_{a},L_{v}](a)+a[L_{a},L_{v}](x).$
Next we show that
$(a,[L_{a},L_{v}](xy))=(a,\left([L_{a},L_{v}](x)\right)y)+(a,\left([L_{a},L_{v}](y)\right)x).$
By Lemma 3.2(2),
${(a,[L_{a},L_{v}](x))=\frac{1}{2}(v,x)-\frac{1}{2}(a,vx_{\frac{1}{2}})=\frac{1}{4}(v,x),}$
so
$\textstyle{(a,[L_{a},L_{v}](xy))=\frac{1}{4}(v,xy)=\frac{1}{4}(v,\frac{1}{2}(a,x)y_{\frac{1}{2}}+\frac{1}{2}(a,y)x_{\frac{1}{2}}+x_{0}y_{\frac{1}{2}}+y_{0}x_{\frac{1}{2}})}$
$\textstyle{=\frac{1}{8}(a,x)(v,y)+\frac{1}{8}(a,y)(v,x)+\frac{1}{4}(v,x_{0}y_{\frac{1}{2}})+\frac{1}{4}(v,y_{0}x_{\frac{1}{2}}).}$
Also, using Lemma 3.2(2),
$\textstyle{(a,\left([L_{a},L_{v}](x)\right)y)=(y,a\big([L_{a},L_{v}](x)\big))=(y,\frac{1}{4}(v,x_{\frac{1}{2}})a-\frac{1}{8}(a,x)v+\frac{1}{4}(vx_{0}))}$
$\textstyle{=\frac{1}{4}(a,y)(v,x)-\frac{1}{8}(a,x)(v,y)+\frac{1}{4}(y_{\frac{1}{2}},vx_{0}).}$
so
$\textstyle{(a,\big([L_{a},L_{v}](x)\big)y)+(a,\big([L_{a},L_{v}](y)\big)x)=\frac{1}{4}(a,y)(v,x)-\frac{1}{8}(a,x)(v,y)+\frac{1}{4}(y_{\frac{1}{2}},vx_{0})}$
$\textstyle{+\frac{1}{4}(a,x)(v,y)-\frac{1}{8}(a,y)(v,x)+\frac{1}{4}(x_{\frac{1}{2}},vy_{0})=}$
$\textstyle{\frac{1}{8}(a,x)(v,y)+\frac{1}{8}(a,y)(v,x)+\frac{1}{4}(x_{\frac{1}{2}},vy_{0})+\frac{1}{4}(y_{\frac{1}{2}},vx_{0}).}$
It remains to show that $\Big([L_{a},L_{b_{\frac{1}{2}}}](xy_{\delta})\Big)_{\lambda}=\Big(\big([L_{a},L_{b_{\frac{1}{2}}}](x)\big)y_{\delta}\Big)_{\lambda}+\Big(\big([L_{a},L_{b_{\frac{1}{2}}}](y_{\delta})\big)x\Big)_{\lambda},$ for $\lambda,\delta\in\{0,\frac{1}{2}\},$ is equivalent to ($i$) (a), (b) and (c). But this follows from Lemma 3.3, 3.4, 3.5 and 3.6. ∎
4. Obtaining identities in AA
We recall Definition 1.1. In particular $a$ and $b$ are axes.
Lemma 4.1.
  1. (1)
    Let $x\in A.$ Then,
    $0=(a,b)ax+(a,b)(b,x)ab-2(a,b)a(bx)+b_{0}x+(b,x)bb_{0}$
    $-2(bx)b_{0}+b_{\frac{1}{2}}x+(b,x)bb_{\frac{1}{2}}-2(bx)b_{\frac{1}{2}}.$
  2. (2)
    $\Big(b_{\frac{1}{2}}^{2}\Big)_{0}=(a,b)\Big[b_{0}^{2}+\Big(b_{\frac{1}{2}}^{2}\Big)_{0}\Big].$
  3. (3)
    1. (a)
      $(b_{\frac{1}{2}},b_{\frac{1}{2}})=(b,b_{\frac{1}{2}})=2(a,b)(1-(a,b)).$
    2. (b)
      $(b,b_{0})=(1-(a,b))^{2}.$
    3. (c)
      $b_{0}^{2}=(1-(a,b))b_{0}$ and $\Big(b_{\frac{1}{2}}^{2}\Big)_{0}=(a,b)b_{0}.$
    4. (d)
      $b_{\frac{1}{2}}^{2}=(a,b)(1-(a,b))a+(a,b)b_{0}.$
    5. (e)
      $b_{\frac{1}{2}}b_{0}=\frac{1}{2}(1-(a,b))b_{\frac{1}{2}}.$
  4. (4)
    $2\Big(b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})\Big)_{0}=x_{0}b_{0}+(b,x_{0})b_{0}-2(x_{0}b_{0})b_{0}.$
  5. (5)
    1. (a)
      $(1-(a,b))\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}-2\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}b_{0}+(b,x_{\frac{1}{2}})b_{0}-2\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}=0.$
    2. (b)
      $\frac{1}{2}(a,b)(1-(a,b))x_{\frac{1}{2}}-2\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}b_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}-2(x_{\frac{1}{2}}b_{0})b_{0}+(1-2(a,b))x_{\frac{1}{2}}b_{0}=0.$
Proof.
(1) Write $b=(a,b)a+b_{0}+b_{\frac{1}{2}}.$ Then, by Lemma 2.4(2) applied to $b,$
$0=b(x+(b,x)b-2bx)=((a,b)a+b_{0}+b_{\frac{1}{2}})(x+(b,x)b-2bx)$
$=(a,b)ax+(a,b)(b,x)ab-2(a,b)a(bx)+b_{0}x+(b,x)bb_{0}$
$-2(bx)b_{0}+b_{\frac{1}{2}}x+(b,x)bb_{\frac{1}{2}}-2(bx)b_{\frac{1}{2}}.$
(2) Replacing $x=a$ in (1) we get
$0=(a,b)a+(a,b)^{2}ab-2(a,b)a(ab)+(a,b)bb_{0}$
$\textstyle{-2(ab)b_{0}+\frac{1}{2}b_{\frac{1}{2}}+(a,b)bb_{\frac{1}{2}}-2(ab)b_{\frac{1}{2}}}$
$=\textstyle{w+\frac{1}{2}(a,b)^{2}b_{\frac{1}{2}}-\frac{1}{2}(a,b)b_{\frac{1}{2}}+(a,b)b_{\frac{1}{2}}b_{0}}$
$\textstyle{-b_{\frac{1}{2}}b_{0}+\frac{1}{2}b_{\frac{1}{2}}+(a,b)b_{0}b_{\frac{1}{2}}+\frac{1}{2}(a,b)^{2}b_{\frac{1}{2}}-(a,b)b_{\frac{1}{2}},}$
where,
$w=(a,b)a+(a,b)^{3}a-2(a,b)^{2}a+(a,b)b_{0}^{2}+(a,b)b_{\frac{1}{2}}^{2}-b_{\frac{1}{2}}^{2}$
$\textstyle{=(a,b)((a,b)-1)^{2}a+\frac{1}{2}((a,b)-1)(b_{\frac{1}{2}},b_{\frac{1}{2}})a}$
$+(a,b)b_{0}^{2}+((a,b)-1)\Big(b_{\frac{1}{2}}^{2}\Big)_{0}.$
Since $w=0,$ we see that $(a,b)(b_{0}^{2}+\Big(b_{\frac{1}{2}}^{2}\Big)_{0})=\Big(b_{\frac{1}{2}}^{2}\Big)_{0}.$
(3) Since $b$ is an idempotent we have
$b$ $=b^{2}=((a,b)a+b_{0}+b_{\frac{1}{2}})^{2}=(a,b)^{2}a+b_{0}^{2}+b_{\frac{1}{2}}^{2}+(a,b)b_{\frac{1}{2}}+2b_{0}b_{\frac{1}{2}}$
$=(a,b)^{2}a+(a,b_{\frac{1}{2}}^{2})a+b_{0}^{2}+\Big(b_{\frac{1}{2}}^{2}\Big)_{0}+(a,b)b_{\frac{1}{2}}+2b_{0}b_{\frac{1}{2}}.$
Since $(a,b_{\frac{1}{2}}^{2})=\frac{1}{2}(b_{\frac{1}{2}},b_{\frac{1}{2}})=\frac{1}{2}(b,b_{\frac{1}{2}}),$ comparing the components of both sides yields (a) and (e) and $b_{0}^{2}+\Big(b_{\frac{1}{2}}^{2}\Big)_{0}=b_{0}.$ Now by (2), $\Big(b_{\frac{1}{2}}^{2}\Big)_{0}=(a,b)b_{0},$ so (c) holds. Since
$1=(b,b)=(a,b)^{2}+(b_{0},b_{0})+(b_{\frac{1}{2}},b_{\frac{1}{2}}),$
(a) implies $(b,b_{0})=(b_{0},b_{0})=(1-(a,b))^{2},$ so (b) holds. Then, since $b_{\frac{1}{2}}^{2}=(a,b_{\frac{1}{2}}^{2})a+\Big(b_{\frac{1}{2}}^{2}\Big)_{0},$ we see that (d) holds.
In the proofs of (4) and (5) ahead, $w$ stands for the $1$-component plus the 0-component of the right side of the equation.
(4) Replacing $x=x_{0}$ in (1) we get,
$0=(a,b)(b,x_{0})ab-2(a,b)a(bx_{0})+b_{0}x_{0}+(b,x_{0})bb_{0}$
$-2(bx_{0})b_{0}+b_{\frac{1}{2}}x_{0}+(b,x_{0})bb_{\frac{1}{2}}-2(bx_{0})b_{\frac{1}{2}}$
$=w+z_{\frac{1}{2}},$
with $w\in\mathbb{F}a+A_{0}(a)$ and $z_{\frac{1}{2}}\in A_{\frac{1}{2}}(a),$ where,
$w=(a,b)^{2}(b,x_{0})a+b_{0}x_{0}+(b,x_{0})b_{0}^{2}-2(x_{0}b_{0})b_{0}+(b,x_{0})b_{\frac{1}{2}}^{2}-2b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})$
$=(a,b)^{2}(b,x_{0})a+x_{0}b_{0}+(b,x_{0})(1-(a,b))b_{0}-2(x_{0}b_{0})b_{0}$
$+(b,x_{0})(a,b)(1-(a,b))a+(b,x_{0})(a,b)b_{0}-(a,b)(b,x_{0})a-2\Big(b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})\Big)_{0}$
$=x_{0}b_{0}+(b,x_{0})b_{0}-2(x_{0}b_{0})b_{0}-2\Big(b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})\Big)_{0},$
where we used that $(a,b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0}))=\frac{1}{2}(b_{\frac{1}{2}}^{2},x_{0})=\frac{1}{2}(a,b)(b,x_{0}),$ and part (2) for $b_{0}^{2}$ and $b_{\frac{1}{2}}^{2}.$ Since $w=0,$ (4) follows.
(5) Replacing $x=x_{\frac{1}{2}}$ in (1) we get,
$\textstyle{0=\frac{1}{2}(a,b)x_{\frac{1}{2}}+(a,b)(b,x_{\frac{1}{2}})ab-2(a,b)a(bx_{\frac{1}{2}})+b_{0}x_{\frac{1}{2}}+(b,x_{\frac{1}{2}})bb_{0}}$
$-2(bx_{\frac{1}{2}})b_{0}+b_{\frac{1}{2}}x_{\frac{1}{2}}+(b,x_{\frac{1}{2}})bb_{\frac{1}{2}}-2(bx_{\frac{1}{2}})b_{\frac{1}{2}}$
$=\textstyle{w+\frac{1}{2}(a,b)x_{\frac{1}{2}}+\frac{1}{2}(a,b)(b,x_{\frac{1}{2}})b_{\frac{1}{2}}-\frac{1}{2}(a,b)^{2}x_{\frac{1}{2}}-(a,b)x_{\frac{1}{2}}b_{0}}$
$+x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})b_{\frac{1}{2}}b_{0}$
$\textstyle{-(a,b)x_{\frac{1}{2}}b_{0}-2(x_{\frac{1}{2}}b_{0})b_{0}+\frac{1}{2}(a,b)(b,x_{\frac{1}{2}})b_{\frac{1}{2}}+(b,x_{\frac{1}{2}})b_{\frac{1}{2}}b_{0}-2(b_{\frac{1}{2}}x_{\frac{1}{2}})b_{\frac{1}{2}}}$
$=\textstyle{w+\frac{1}{2}(a,b)(1-(a,b))x_{\frac{1}{2}}+(a,b)(b,x_{\frac{1}{2}})b_{\frac{1}{2}}+(1-2(a,b))x_{\frac{1}{2}}b_{0}}$
$+(1-(a,b))(b,x_{\frac{1}{2}})b_{\frac{1}{2}}-2(x_{\frac{1}{2}}b_{0})b_{0}-2(b_{\frac{1}{2}}x_{\frac{1}{2}})b_{\frac{1}{2}}$
$\textstyle{=w+\frac{1}{2}(a,b)(1-(a,b))x_{\frac{1}{2}}+(b,x_{\frac{1}{2}})b_{\frac{1}{2}}+(1-2(a,b))x_{\frac{1}{2}}b_{0}}$
$-2(x_{\frac{1}{2}}b_{0})b_{0}-2(b_{\frac{1}{2}}x_{\frac{1}{2}})b_{\frac{1}{2}}$
$\textstyle{=w+\frac{1}{2}(a,b)(1-(a,b))x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}+(1-2(a,b))x_{\frac{1}{2}}b_{0}}$
$-2(x_{\frac{1}{2}}b_{0})b_{0}-2\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}b_{\frac{1}{2}},$
with $w\in\mathbb{F}a+A_{0}(a),$ where we used $b_{\frac{1}{2}}x_{\frac{1}{2}}=\frac{1}{2}(b,x_{\frac{1}{2}})a+\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}.$ This shows part (b). Now
$w=(a,b)^{2}(b,x_{\frac{1}{2}})a-2(a,b)a(b_{\frac{1}{2}}x_{\frac{1}{2}})+(b,x_{\frac{1}{2}})b_{0}^{2}-2(b_{\frac{1}{2}}x_{\frac{1}{2}})b_{0}+$
$b_{\frac{1}{2}}x_{\frac{1}{2}}+(b,x_{\frac{1}{2}})b_{\frac{1}{2}}^{2}-(a,b)x_{\frac{1}{2}}b_{\frac{1}{2}}-2(b_{0}x_{\frac{1}{2}})b_{\frac{1}{2}}$
$\textstyle{=(b,x_{\frac{1}{2}})((a,b)-1)((a,b)-\frac{1}{2})a+\frac{1}{2}(b_{\frac{1}{2}},b_{\frac{1}{2}})(b,x_{\frac{1}{2}})a-(b_{\frac{1}{2}},b_{0}x_{\frac{1}{2}})a}$
$+(b,x_{\frac{1}{2}})b_{0}^{2}-2(b_{\frac{1}{2}}x_{\frac{1}{2}})b_{0}+(1-(a,b))\Big(x_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}$
$+(b,x_{\frac{1}{2}})\Big(b_{\frac{1}{2}}^{2}\Big)_{0}-2\Big((b_{0}x_{\frac{1}{2}})b_{\frac{1}{2}}\Big)_{0},$
where the last equality follows from
$b_{\frac{1}{2}}x_{\frac{1}{2}}=\frac{1}{2}(b,x_{\frac{1}{2}})a+\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}\text{ so}$
$(a,b)^{2}(b,x_{\frac{1}{2}})a-2(a,b)a(b_{\frac{1}{2}}x_{\frac{1}{2}})+(1-(a,b))b_{\frac{1}{2}}x_{\frac{1}{2}}=$
$\textstyle{\Big[(a,b)^{2}(b,x_{\frac{1}{2}})-(a,b)(b,x_{\frac{1}{2}})+\frac{1}{2}(1-(a,b))(b,x_{\frac{1}{2}})\Big]a+(1-(a,b))\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}=}$
$\textstyle{\Big[(b,x_{\frac{1}{2}})(a,b)((a,b)-1))+\frac{1}{2}(1-(a,b))(b,x_{\frac{1}{2}})\Big]a+(1-(a,b))\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}=}$
$\textstyle{(b,x_{\frac{1}{2}})((a,b)-1)((a,b)-\frac{1}{2})a+(1-(a,b))\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}.}$
We have
$0=(b,x_{\frac{1}{2}})b_{0}^{2}-2(b_{\frac{1}{2}}x_{\frac{1}{2}})b_{0}+(1-(a,b))\Big(x_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}$
$+(b,x_{\frac{1}{2}})\Big(b_{\frac{1}{2}}^{2}\Big)_{0}-2\Big((b_{0}x_{\frac{1}{2}})b_{\frac{1}{2}}\Big)_{0}$
$=(b,x_{\frac{1}{2}})(1-(a,b))b_{0}-2\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}b_{0}+(1-(a,b))\Big(x_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}$
$+(b,x_{\frac{1}{2}})(a,b)b_{0}-2\Big((b_{0}x_{\frac{1}{2}})b_{\frac{1}{2}}\Big)_{0}$
$=(b,x_{\frac{1}{2}})b_{0}-2\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}b_{0}+(1-(a,b))\Big(x_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}-2\Big((b_{0}x_{\frac{1}{2}})b_{\frac{1}{2}}\Big)_{0},$
so part (a) holds. ∎
Lemma 4.2.
Let $a,b$ be axes in $A.$ Then
  1. (1)
    a(x(y+(a,y)a2ay))=(ax)(y+(a,y)a2ay),a(x(y+(a,y)a-2ay))=(ax)(y+(a,y)a-2ay), for all $x,y\in A,$ that is a(xy)2a((ay)x)=(ax)y2(ax)(ay).a(xy)-2a((ay)x)=(ax)y-2(ax)(ay).
  2. (2)
    For all $x,y\in A,$
    $(a,b)(1-(a,b))a(xy)+b_{0}(xy)+b_{\frac{1}{2}}(xy)-2\Big((a,b)b_{0}((ay)x)+(a,b)b_{\frac{1}{2}}((ay)x)\Big)$
    $-2\Big((a,b)a((b_{0}y)x)+b_{0}((b_{0}y)x)+b_{\frac{1}{2}}((b_{0}y)x)$
    $+(a,b)a((b_{\frac{1}{2}}y)x)+b_{0}((b_{\frac{1}{2}}y)x)+b_{\frac{1}{2}}((b_{\frac{1}{2}}y)x)\Big)$
    $(a,b)(1-(a,b))(ax)y+(b_{0}x)y+(b_{\frac{1}{2}}x)y-2\Big((a,b)(ax)(b_{0}y)+(a,b)(ax)(b_{\frac{1}{2}}y)\Big)$
    $-2\Big((a,b)(b_{0}x)(ay)+(b_{0}x)(b_{0}y)+(b_{0}x)(b_{\frac{1}{2}}y)\Big)$
    $-2\Big((a,b)(b_{\frac{1}{2}}x)(ay)+(b_{\frac{1}{2}}x)(b_{0}y)+(b_{\frac{1}{2}}x)(b_{\frac{1}{2}}y)\Big).$
  3. (3)
    Taking $x=x_{0}\in A_{0}(a)$ and $y=a$ in (2) yields
    $(1-(a,b))b_{\frac{1}{2}}x_{0}-2b_{\frac{1}{2}}(x_{0}b_{0})+2(b_{\frac{1}{2}}x_{0})b_{0}=0.$
  4. (4)
    Taking $x=a$ and $y=y_{\frac{1}{2}}$ in (2), we get
    1. (a)
      $\frac{1}{2}(a,b)(1-(a,b))y_{\frac{1}{2}}-2b_{\frac{1}{2}}\Big(y_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}+\frac{1}{2}(b,y_{\frac{1}{2}})b_{\frac{1}{2}}+2(y_{\frac{1}{2}}b_{0})b_{0}-y_{\frac{1}{2}}b_{0}=0.$
    2. (b)
      $(x_{\frac{1}{2}}b_{0})b_{0}=\frac{1}{2}(1-(a,b))x_{\frac{1}{2}}b_{0}.$
    3. (c)
      $b_{\frac{1}{2}}\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}=\frac{1}{4}(a,b)(1-(a,b))x_{\frac{1}{2}}-\frac{1}{2}(a,b)x_{\frac{1}{2}}b_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}.$
    4. (d)
      $b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{\frac{1}{2}})=\frac{1}{4}(a,b)(1-(a,b))x_{\frac{1}{2}}-\frac{1}{2}(a,b)x_{\frac{1}{2}}b_{0}+\frac{1}{2}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}.$
  5. (5)
    Taking $y=a$ and $x=x_{\frac{1}{2}}$ in (2) we get
    1. (a)
      $(1-(a,b))\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}-2(b_{\frac{1}{2}}x_{\frac{1}{2}})b_{0}+2\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}=0.$
    2. (b)
      $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}=\frac{1}{4}(b,x_{\frac{1}{2}})b_{0}.$
    3. (c)
      $b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})=\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}})a+\frac{1}{4}(b,x_{\frac{1}{2}})b_{0}.$
    4. (d)
      From (a) and (b) it follows:
      ${(b_{\frac{1}{2}}x_{\frac{1}{2}})b_{0}=\frac{1}{2}(1-(a,b))\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})b_{0}.}$
  6. (6)
    Taking $x=x_{0},$ and $y=y_{0}$ in (2) we get
    $b_{0}(x_{0}y_{0})+b_{\frac{1}{2}}(x_{0}y_{0})-2\Big[b_{0}((b_{0}y_{0})x_{0})+b_{\frac{1}{2}}((b_{0}y_{0})x_{0})$
    $\textstyle{+\frac{1}{2}(a,b)(b_{\frac{1}{2}}y_{0})x_{0}+b_{0}((b_{\frac{1}{2}}y_{0})x_{0})+b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{0})x_{0})\Big]}$
    $=(b_{0}x_{0})y_{0}+(b_{\frac{1}{2}}x_{0})y_{0}$
    $-2\Big[(b_{0}x_{0})(b_{0}y_{0})+(b_{0}x_{0})(b_{\frac{1}{2}}y_{0})\Big]$
    $-2\Big[(b_{\frac{1}{2}}x_{0})(b_{0}y_{0})+(b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{0})\Big].$
  7. (7)
    For all $x_{0},y_{0}\in A_{0}(a),$
    $b_{\frac{1}{2}}(x_{0}y_{0})+2(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})+2(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})$
    $=(b_{\frac{1}{2}}x_{0})y_{0}+2b_{\frac{1}{2}}(x_{0}(y_{0}b_{0}))+(a,b)(b_{\frac{1}{2}}y_{0})x_{0}+2((b_{\frac{1}{2}}y_{0})x_{0})b_{0}.$
  8. (8)
    Interchanging $x_{0}$ and $y_{0}$ in (7) we get
    1. (a)
      By subtracting we get
      $(1-(a,b))(b_{\frac{1}{2}}x_{0})y_{0}-(1-(a,b))(b_{\frac{1}{2}}y_{0})x_{0}$
      $=2b_{\frac{1}{2}}(y_{0}(x_{0}b_{0}))-2b_{\frac{1}{2}}(x_{0}(y_{0}b_{0}))+2((b_{\frac{1}{2}}x_{0})y_{0})b_{0}-2((b_{\frac{1}{2}}y_{0})x_{0})b_{0}.$
    2. (b)
      By adding we get
      $2b_{\frac{1}{2}}(x_{0}y_{0})+4(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})+4(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})$
      $=(1+(a,b))(b_{\frac{1}{2}}x_{0})y_{0}+(1+(a,b))(b_{\frac{1}{2}}y_{0})x_{0}+2b_{\frac{1}{2}}((y_{0}b_{0})x_{0})+2b_{\frac{1}{2}}((x_{0}b_{0})y_{0})$
      $+2((b_{\frac{1}{2}}y_{0})x_{0})b_{0}+2((b_{\frac{1}{2}}x_{0})y_{0})b_{0}.$
  9. (9)
    For all $x_{0},y_{0}\in A_{0}(a),$
    $(x_{0}y_{0})b_{0}+2(x_{0}b_{0})(y_{0}b_{0})+2\Big((b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}$
    $=y_{0}(x_{0}b_{0})+2(x_{0}(y_{0}b_{0}))b_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{0})x_{0})\Big)_{0}.$
  10. (10)
    Interchanging $x_{0}$ and $y_{0}$ in (9), we get
    1. (a)
      by subtracting,
      $2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}x_{0})y_{0})\Big)_{0}-2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{0})x_{0})\Big)_{0}$
      $=y_{0}(x_{0}b_{0})+2(x_{0}(y_{0}b_{0}))b_{0}-x_{0}(y_{0}b_{0})-2(y_{0}(x_{0}b_{0}))b_{0}.$
    2. (b)
      By adding,
      $4(x_{0}b_{0})(y_{0}b_{0})+4\Big((b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}-y_{0}(x_{0}b_{0})-x_{0}(y_{0}b_{0})+(x_{0}y_{0})b_{0}$
      $=2(x_{0}(y_{0}b_{0}))b_{0}+2(y_{0}(x_{0}b_{0}))b_{0}$
      $+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{0})x_{0})\Big)_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}x_{0})y_{0})\Big)_{0}-(x_{0}y_{0})b_{0}.$
Proof.
(1) Follows from Seress’ lemma 2.2, and Lemma 2.4(1).
(2) By (1),
$b(xy)-2b((by)x)=(bx)y-2(bx)(by)\quad\iff$
$((a,b)a+b_{0}+b_{\frac{1}{2}})(xy)-2((a,b)a+b_{0}+b_{\frac{1}{2}})((((a,b)a+b_{0}+b_{\frac{1}{2}})y)x)$
$=(((a,b)a+b_{0}+b_{\frac{1}{2}})x)y-2(((a,b)a+b_{0}+b_{\frac{1}{2}})x)(((a,b)a+b_{0}+b_{\frac{1}{2}})y)$
$\iff$
$(a,b)a(xy)+b_{0}(xy)+b_{\frac{1}{2}}(xy)-2\Big[(a,b)^{2}a((ay)x))+(a,b)b_{0}((ay)x)+(a,b)b_{\frac{1}{2}}((ay)x)\Big]$
$-2\Big[(a,b)a((b_{0}y)x)+b_{0}((b_{0}y)x)+b_{\frac{1}{2}}((b_{0}y)x)+(a,b)a((b_{\frac{1}{2}}y)x)+b_{0}((b_{\frac{1}{2}}y)x)+b_{\frac{1}{2}}((b_{\frac{1}{2}}y)x)\Big]$
$=(a,b)(ax)y+(b_{0}x)y+(b_{\frac{1}{2}}x)y-2\Big[(a,b)^{2}(ax)(ay)+(a,b)(ax)(b_{0}y)+(a,b)(ax)(b_{\frac{1}{2}}y)\Big]$
$-2\Big[(a,b)(b_{0}x)(ay)+(b_{0}x)(b_{0}y)+(b_{0}x)(b_{\frac{1}{2}}y)\Big]$
$-2\Big[(a,b)(b_{\frac{1}{2}}x)(ay)+(b_{\frac{1}{2}}x)(b_{0}y)+(b_{\frac{1}{2}}x)(b_{\frac{1}{2}}y)\Big]$
$\iff$
$(a,b)(1-(a,b))a(xy)+b_{0}(xy)+b_{\frac{1}{2}}(xy)-2\Big[(a,b)b_{0}((ay)x)+(a,b)b_{\frac{1}{2}}((ay)x)\Big]$
$-2\Big[(a,b)a((b_{0}y)x)+b_{0}((b_{0}y)x)+b_{\frac{1}{2}}((b_{0}y)x)+(a,b)a((b_{\frac{1}{2}}y)x)+b_{0}((b_{\frac{1}{2}}y)x)+b_{\frac{1}{2}}((b_{\frac{1}{2}}y)x)\Big]$
$=(a,b)(1-(a,b))(ax)y+(b_{0}x)y+(b_{\frac{1}{2}}x)y-2\Big[(a,b)(ax)(b_{0}y)+(a,b)(ax)(b_{\frac{1}{2}}y)\Big]$
$-2\Big[(a,b)(b_{0}x)(ay)+(b_{0}x)(b_{0}y)+(b_{0}x)(b_{\frac{1}{2}}y)\Big]$
$-2\Big[(a,b)(b_{\frac{1}{2}}x)(ay)+(b_{\frac{1}{2}}x)(b_{0}y)+(b_{\frac{1}{2}}x)(b_{\frac{1}{2}}y)\Big]$
where we used $(a,b)^{2}\Big[a(xy)-2a((ay)x)\Big]=(a,b)^{2}\Big[(ax)y-2(ax)(ay)\Big].$
(3) Taking $y=a$ and $x=x_{0}$ in (2) we get,
$\textstyle{-\frac{1}{2}(a,b)(b_{\frac{1}{2}}x_{0})-b_{0}(b_{\frac{1}{2}}x_{0})-b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})}$
$\textstyle{=\frac{1}{2}b_{\frac{1}{2}}x_{0}-(b_{0}x_{0})b_{\frac{1}{2}}-(a,b)(b_{\frac{1}{2}}x_{0})-(b_{\frac{1}{2}}x_{0})b_{\frac{1}{2}}\iff}$
$\textstyle{\frac{1}{2}((a,b)-1)b_{\frac{1}{2}}x_{0}-b_{0}(b_{\frac{1}{2}}x_{0})=-(b_{0}x_{0})b_{\frac{1}{2}}.}$
This shows (3).
(4) (a): Replacing $x=a$ and $y=y_{\frac{1}{2}}$ in (2) we get
$\textstyle{\frac{1}{4}(a,b)(1-(a,b))y_{\frac{1}{2}}+\frac{1}{2}y_{\frac{1}{2}}b_{0}+\frac{1}{2}y_{\frac{1}{2}}b_{\frac{1}{2}}-2\Big[\frac{1}{4}(a,b)y_{\frac{1}{2}}b_{0}+\frac{1}{4}(a,b)y_{\frac{1}{2}}b_{\frac{1}{2}}\Big]}$
$\textstyle{-2\Big[\frac{1}{4}(a,b)y_{\frac{1}{2}}b_{0}+\frac{1}{2}(y_{\frac{1}{2}}b_{0})b_{0}+\frac{1}{2}(y_{\frac{1}{2}}b_{0})b_{\frac{1}{2}}+(a,b)(a,b_{\frac{1}{2}}y_{\frac{1}{2}})a+\frac{1}{2}(a,b_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}\Big]}$
$\textstyle{=\frac{1}{2}(a,b)(1-(a,b))y_{\frac{1}{2}}+\frac{1}{2}b_{\frac{1}{2}}y_{\frac{1}{2}}-2\Big[\frac{1}{2}(a,b)y_{\frac{1}{2}}b_{0}+(a,b)(a,b_{\frac{1}{2}}y_{\frac{1}{2}})a\Big]}$
$\textstyle{-2\Big[\frac{1}{4}(a,b)y_{\frac{1}{2}}b_{\frac{1}{2}}+\frac{1}{2}(y_{\frac{1}{2}}b_{0})b_{\frac{1}{2}}+\frac{1}{2}b_{\frac{1}{2}}(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big].}$
Hence, comparing the $\frac{1}{2}$-components, we get
$\textstyle{\frac{1}{4}(a,b)(1-(a,b))y_{\frac{1}{2}}+\frac{1}{2}y_{\frac{1}{2}}b_{0}-\frac{1}{2}(a,b)y_{\frac{1}{2}}b_{0}}$
$\textstyle{-\frac{1}{2}(a,b)y_{\frac{1}{2}}b_{0}-(y_{\frac{1}{2}}b_{0})b_{0}-(a,y_{\frac{1}{2}}b_{\frac{1}{2}})b_{\frac{1}{2}}}$
$\textstyle{=\frac{1}{2}(a,b)(1-(a,b))y_{\frac{1}{2}}-(a,b)y_{\frac{1}{2}}b_{0}}$
$\textstyle{-\frac{1}{2}(a,b_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-\Big(y_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}b_{\frac{1}{2}}.}$
Hence
$\textstyle{\frac{1}{2}y_{\frac{1}{2}}b_{0}-(y_{\frac{1}{2}}b_{0})b_{0}=\frac{1}{4}(a,b)(1-(a,b))y_{\frac{1}{2}}+\frac{1}{4}(b,y_{\frac{1}{2}})b_{\frac{1}{2}}-b_{\frac{1}{2}}\Big(b_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0},}$
or
${\frac{1}{2}(a,b)(1-(a,b))y_{\frac{1}{2}}-2b_{\frac{1}{2}}\Big(b_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}+\frac{1}{2}(b,y_{\frac{1}{2}})b_{\frac{1}{2}}+2(y_{\frac{1}{2}}b_{0})b_{0}-y_{\frac{1}{2}}b_{0}=0.}$
(b): Comparing (a) with Lemma 4.1(5)(b) which states:
$\textstyle{\frac{1}{2}(a,b)(1-(a,b))x_{\frac{1}{2}}-2b_{\frac{1}{2}}\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}+\frac{1}{2}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}}$
$-2(x_{\frac{1}{2}}b_{0})b_{0}+(1-2(a,b))x_{\frac{1}{2}}b_{0}=0,$
we get
$2(x_{\frac{1}{2}}b_{0})b_{0}-x_{\frac{1}{2}}b_{0}=-2(x_{\frac{1}{2}}b_{0})b_{0}+(1-2(a,b))x_{\frac{1}{2}}b_{0},$
which proves (b).
(c) By lemma 4.1(5)(b), and by (b),
$\textstyle{\frac{1}{2}(a,b)(1-(a,b))x_{\frac{1}{2}}-2b_{\frac{1}{2}}\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}+\frac{1}{2}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}}$
$-2(x_{\frac{1}{2}}b_{0})b_{0}+(1-2(a,b))x_{\frac{1}{2}}b_{0}=0$
$\iff$
$\textstyle{\frac{1}{2}(a,b)(1-(a,b))x_{\frac{1}{2}}-2\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}b_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}-(a,b)x_{\frac{1}{2}}b_{0}=0.}$
(d) Note that $(a,b_{\frac{1}{2}}x_{\frac{1}{2}})=\frac{1}{2}(b,x_{\frac{1}{2}}).$ Hence, by (c),
$b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{\frac{1}{2}})=b_{\frac{1}{2}}\Big[(a,b_{\frac{1}{2}}x_{\frac{1}{2}})a+\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}\Big]$
$\textstyle{=\frac{1}{4}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{4}(a,b)(1-(a,b))x_{\frac{1}{2}}+\frac{1}{4}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}-\frac{1}{2}(a,b)x_{\frac{1}{2}}b_{0}.}$
(5) Taking $y=a$ and $x=x_{\frac{1}{2}}$ in (2) we get
$\textstyle{\frac{1}{4}(a,b)(1-(a,b))x_{\frac{1}{2}}+\frac{1}{2}x_{\frac{1}{2}}b_{0}+\frac{1}{2}b_{\frac{1}{2}}x_{\frac{1}{2}}-(a,b)x_{\frac{1}{2}}b_{0}-(a,b)b_{\frac{1}{2}}x_{\frac{1}{2}}}$
$\textstyle{-(a,b)a(b_{\frac{1}{2}}x_{\frac{1}{2}})-b_{0}(b_{\frac{1}{2}}x_{\frac{1}{2}})-b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{\frac{1}{2}})\quad=}$
$\textstyle{\frac{1}{4}(a,b)(1-(a,b))x_{\frac{1}{2}}+\frac{1}{2}x_{\frac{1}{2}}b_{0}+a(b_{\frac{1}{2}}x_{\frac{1}{2}})-\frac{1}{2}(a,b)x_{\frac{1}{2}}b_{\frac{1}{2}}}$
$-(a,b)x_{\frac{1}{2}}b_{0}-(x_{\frac{1}{2}}b_{0})b_{\frac{1}{2}}-2(a,b)a(b_{\frac{1}{2}}x_{\frac{1}{2}})-(b_{\frac{1}{2}}x_{\frac{1}{2}})b_{\frac{1}{2}}$
$\iff$
$\textstyle{\frac{1}{2}b_{\frac{1}{2}}x_{\frac{1}{2}}-\frac{1}{2}(a,b)b_{\frac{1}{2}}x_{\frac{1}{2}}-(a,b)a(b_{\frac{1}{2}}x_{\frac{1}{2}})-b_{0}(b_{\frac{1}{2}}x_{\frac{1}{2}})}$
$\textstyle{=a(b_{\frac{1}{2}}x_{\frac{1}{2}})-(b_{0}x_{\frac{1}{2}})b_{\frac{1}{2}}-2(a,b)a(b_{\frac{1}{2}}x_{\frac{1}{2}})}\quad\iff$
$\textstyle{\frac{1}{2}(1-(a,b))b_{\frac{1}{2}}x_{\frac{1}{2}}+((a,b)-1)a(b_{\frac{1}{2}}x_{\frac{1}{2}})+(b_{0}x_{\frac{1}{2}})b_{\frac{1}{2}}-b_{0}(b_{\frac{1}{2}}x_{\frac{1}{2}})=0.}$
Writing $b_{\frac{1}{2}}x_{\frac{1}{2}}=\frac{1}{2}(b,x_{\frac{1}{2}})a+\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0},$ and $b_{\frac{1}{2}}(b_{0}x_{\frac{1}{2}})=\frac{1}{2}(b,b_{0}x_{\frac{1}{2}})a+\Big(b_{\frac{1}{2}}(b_{0}x_{\frac{1}{2}})\Big)_{0},$ we get $((a,b)-1)a(b_{\frac{1}{2}}x_{\frac{1}{2}})=\frac{1}{2}((a,b)-1)(b,x_{\frac{1}{2}})a,$ so
$\textstyle{\frac{1}{4}((a,b)-1)(b,x_{\frac{1}{2}})a+\frac{1}{2}(b,b_{0}x_{\frac{1}{2}})a}$
$\textstyle{+\frac{1}{2}(1-(a,b))\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}+\Big((b_{0}x_{\frac{1}{2}})b_{\frac{1}{2}}\Big)_{0}-b_{0}(b_{\frac{1}{2}}x_{\frac{1}{2}})=0.}$
Now $\frac{1}{2}(b,b_{0}x_{\frac{1}{2}})=\frac{1}{2}(b_{\frac{1}{2}}b_{0},x_{\frac{1}{2}})=\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}}),$ so part (a) holds. Then, (b) follows from (a) and Lemma 4.1(5)(a).
(c) We only need to compute $(a,b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))=\frac{1}{2}(b_{\frac{1}{2}}b_{0},x_{\frac{1}{2}})=\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}}).$ (7) Comparing the $\frac{1}{2}$-components in (6) we get
$\textstyle{b_{\frac{1}{2}}(x_{0}y_{0})-2\Big[b_{\frac{1}{2}}((y_{0}b_{0})x_{0})+\frac{1}{2}(a,b)(b_{\frac{1}{2}}y_{0})x_{0}+((b_{\frac{1}{2}}y_{0})x_{0})b_{0}\Big]}$
$=(b_{\frac{1}{2}}x_{0})y_{0}-2(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})-2(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})$
$\iff$
$b_{\frac{1}{2}}(x_{0}y_{0})+2(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})+2(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})$
$=(b_{\frac{1}{2}}x_{0})y_{0}+2b_{\frac{1}{2}}((y_{0}b_{0})x_{0})+(a,b)(b_{\frac{1}{2}}y_{0})x_{0}+2((b_{\frac{1}{2}}y_{0})x_{0})b_{0}.$
(8) Interchanging $x_{0}$ and $y_{0}$ in (7) we get, by subtracting,
$(b_{\frac{1}{2}}x_{0})y_{0}+2b_{\frac{1}{2}}((y_{0}b_{0})x_{0})+(a,b)(b_{\frac{1}{2}}y_{0})x_{0}+2((b_{\frac{1}{2}}y_{0})x_{0})b_{0}$
$=(b_{\frac{1}{2}}y_{0})x_{0}+2b_{\frac{1}{2}}((x_{0}b_{0})y_{0})+(a,b)(b_{\frac{1}{2}}x_{0})y_{0}+2((b_{\frac{1}{2}}x_{0})y_{0})b_{0}$
$\iff$
$(1-(a,b))(b_{\frac{1}{2}}x_{0})y_{0}-(1-(a,b))(b_{\frac{1}{2}}y_{0})x_{0}$
$=2b_{\frac{1}{2}}((x_{0}b_{0})y_{0})-2b_{\frac{1}{2}}((y_{0}b_{0})x_{0})+2((b_{\frac{1}{2}}x_{0})y_{0})b_{0}-2((b_{\frac{1}{2}}y_{0})x_{0})b_{0}.$
The other part of (8) is clear.
(9) Comparing the 0-components in (6) we get,
$b_{0}(x_{0}y_{0})-2b_{0}((b_{0}y_{0})x_{0})-2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{0})x_{0})\Big)_{0}$
$=(b_{0}x_{0})y_{0}-2(b_{0}x_{0})(b_{0}y_{0})-2\Big((b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}$
$\iff$
$b_{0}(x_{0}y_{0})+2(b_{0}x_{0})(b_{0}y_{0})+2\Big((b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}$
$=(b_{0}x_{0})y_{0}+2b_{0}((b_{0}y_{0})x_{0})+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{0})x_{0})\Big)_{0}.$
(10) This is immediate from (9). ∎
Lemma 4.3.
Let $x_{\frac{1}{2}}\in A_{\frac{1}{2}}(a)$ and $y_{0}\in A_{0}(a),$ then
  1. (1)
    $(x_{\frac{1}{2}}y_{0})b_{0}+b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})$
    $-2\Big[(x_{\frac{1}{2}}(y_{0}b_{0}))b_{0}+b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))+(a,b)a((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})+(x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0}))b_{0}$
    $+b_{\frac{1}{2}}(x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0}))\Big]\quad=$
    $(x_{\frac{1}{2}}b_{0})y_{0}+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-(a,b)x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0})$
    $-2\Big[(x_{\frac{1}{2}}b_{0})(y_{0}b_{0})+(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big]$
    $-2\Big[(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})+(b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{0})\Big].$
  2. (2)
    $(x_{\frac{1}{2}}y_{0})b_{0}-2(x_{\frac{1}{2}}(y_{0}b_{0}))b_{0}-2b_{\frac{1}{2}}(x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0}))\quad=$
    $(x_{\frac{1}{2}}b_{0})y_{0}-2(x_{\frac{1}{2}}b_{0})(y_{0}b_{0})-2(b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{0}).$
  3. (3)
    $b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})+(a,b)(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+2(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})$
    $+2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-2(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))-2(a,b)a(x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0}))-2((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}$
    $=0.$
Proof.
(1) taking $x=x_{\frac{1}{2}}$ and $y=y_{0}$ in Lemma 4.2(2), we get
$(a,b)(1-(a,b))a(x_{\frac{1}{2}}y_{0})+b_{0}(x_{\frac{1}{2}}y_{0})+b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})-2\Big[(a,b)b_{0}((ay_{0})x_{\frac{1}{2}})+(a,b)b_{\frac{1}{2}}((ay_{0})x_{\frac{1}{2}})\Big]$
$-2\Big[(a,b)a((b_{0}y_{0})x_{\frac{1}{2}})+b_{0}((b_{0}y_{0})x_{\frac{1}{2}})+b_{\frac{1}{2}}((b_{0}y_{0})x_{\frac{1}{2}})+(a,b)a((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})+b_{0}((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})$
$+b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})\Big]$
$(a,b)(1-(a,b))(ax_{\frac{1}{2}})y_{0}+(b_{0}x_{\frac{1}{2}})y_{0}+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-2\Big[(a,b)(ax_{\frac{1}{2}})(b_{0}y_{0})+(a,b)(ax_{\frac{1}{2}})(b_{\frac{1}{2}}y_{0})\Big]$
$-2\Big[(a,b)(b_{0}x_{\frac{1}{2}})(ay_{0})+(b_{0}x_{\frac{1}{2}})(b_{0}y_{0})+(b_{0}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{0})\Big]$
$-2\Big[(a,b)(b_{\frac{1}{2}}x_{\frac{1}{2}})(ay_{0})+(b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{0}y_{0})+(b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{0})\Big].$
So,
$\textstyle{\frac{1}{2}(a,b)(1-(a,b))x_{\frac{1}{2}}y_{0}+(x_{\frac{1}{2}}y_{0})b_{0}+b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})}$
$\textstyle{-2\Big[\frac{1}{2}(a,b)x_{\frac{1}{2}}(y_{0}b_{0})+(x_{\frac{1}{2}}(y_{0}b_{0}))b_{0}+b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))+(a,b)a((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})+((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}}$
$+b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})\Big]$
$\textstyle{\frac{1}{2}(a,b)(1-(a,b))x_{\frac{1}{2}}y_{0}+(x_{\frac{1}{2}}b_{0})y_{0}+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-2\Big[\frac{1}{2}(a,b)x_{\frac{1}{2}}(y_{0}b_{0})+\frac{1}{2}(a,b)x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0})\Big]}$
$-2\Big[(x_{\frac{1}{2}}b_{0})(y_{0}b_{0})+(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big)$
$-2\Big((b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})+(b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{0})\Big].$
Hence,
$(x_{\frac{1}{2}}y_{0})b_{0}+b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0}))$
$-2\Big[(x_{\frac{1}{2}}(y_{0}b_{0}))b_{0}+b_{\frac{1}{2}}(x_{\frac{1}{2}}(b_{0}y_{0}))+(a,b)a((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})+((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}$
$+b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})\Big]$
$=\textstyle{(x_{\frac{1}{2}}b_{0})y_{0}+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-2\Big(\frac{1}{2}(a,b)x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0})\Big)}$
$-2\Big[(x_{\frac{1}{2}}b_{0})(y_{0}b_{0})+(b_{0}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{0})\Big]$
$-2\Big[(b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{0}y_{0})+(b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{0})\Big].$
(2) Comparing the $\frac{1}{2}$-components in (1) we get
$(x_{\frac{1}{2}}y_{0})b_{0}-2(x_{\frac{1}{2}}(y_{0}b_{0}))b_{0}-2b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})$
$=(x_{\frac{1}{2}}b_{0})y_{0}-2(x_{\frac{1}{2}}b_{0})(y_{0}b_{0})-2(b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{0}).$
This shows (2).
(3) Comparing the $1$-component plus the 0-components on both sides of (1) gives:
$b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})-2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))-2(a,b)a(x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0}))-2(x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0}))b_{0}$
$=(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-(a,b)x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0})-2(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})-2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0}).\qed$
Lemma 4.4.
Let $x_{0}\in A_{0}(a)$ and $y_{\frac{1}{2}}\in A_{\frac{1}{2}}(a),then$
  1. (1)
    $\textstyle{\frac{1}{2}(a,b)(1-(a,b))y_{\frac{1}{2}}x_{0}+(1-(a,b))(y_{\frac{1}{2}}x_{0})b_{0}+(1-(a,b))b_{\frac{1}{2}}(y_{\frac{1}{2}}x_{0})}$
    $-(a,b)(y_{\frac{1}{2}}b_{0})x_{0}-2((y_{\frac{1}{2}}b_{0})x_{0})b_{0}-2b_{\frac{1}{2}}((y_{\frac{1}{2}}b_{0})x_{0})$
    $-2((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{0})b_{0}-2b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{0})$
    $=(1-(a,b))y_{\frac{1}{2}}(x_{0}b_{0})+(1-(a,b))y_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})$
    $-2(y_{\frac{1}{2}}b_{0})(b_{0}x_{0})-2(b_{\frac{1}{2}}y_{\frac{1}{2}})(b_{0}x_{0})$
    $-2(b_{\frac{1}{2}}x_{0})(y_{\frac{1}{2}}b_{0})-2(b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}}).$
  2. (2)
    $\textstyle{\frac{1}{2}(a,b)(1-(a,b))y_{\frac{1}{2}}x_{0}+(1-(a,b))(y_{\frac{1}{2}}x_{0})b_{0}-(a,b)(y_{\frac{1}{2}}b_{0})x_{0}-2((y_{\frac{1}{2}}b_{0})x_{0})b_{0}}$
    $-2b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{0})$
    $=(1-(a,b))y_{\frac{1}{2}}(x_{0}b_{0})-2(y_{\frac{1}{2}}b_{0})(b_{0}x_{0})-2(b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}}).$
  3. (3)
    $(1-(a,b))b_{\frac{1}{2}}(y_{\frac{1}{2}}x_{0})-(1-(a,b))(b_{\frac{1}{2}}x_{0})y_{\frac{1}{2}}+2(b_{\frac{1}{2}}x_{0})(y_{\frac{1}{2}}b_{0})$
    $+2(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{0}b_{0})-2((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{0})b_{0}-2b_{\frac{1}{2}}((y_{\frac{1}{2}}b_{0})x_{0})$
    $=0.$
Proof.
(1) Replacing $x=x_{0}$ and $y=y_{\frac{1}{2}}$ in Lemma 4.2(2) we get
$(a,b)(1-(a,b))a(x_{0}y_{\frac{1}{2}})+b_{0}(x_{0}y_{\frac{1}{2}})+b_{\frac{1}{2}}(x_{0}y_{\frac{1}{2}})-2\Big((a,b)b_{0}((ay_{\frac{1}{2}})x_{0})+(a,b)b_{\frac{1}{2}}((ay_{\frac{1}{2}})x_{0})\Big)$
$-2\Big((a,b)a((b_{0}y_{\frac{1}{2}})x_{0})+b_{0}((b_{0}y_{\frac{1}{2}})x_{0})+b_{\frac{1}{2}}((b_{0}y_{\frac{1}{2}})x_{0})+(a,b)a((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{0})$
$+b_{0}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{0})+b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{0})\Big)$
$(a,b)(1-(a,b))(ax_{0})y_{\frac{1}{2}}+(b_{0}x_{0})y_{\frac{1}{2}}+(b_{\frac{1}{2}}x_{0})y_{\frac{1}{2}}$
$-2\Big((a,b)(ax_{0})(b_{0}y_{\frac{1}{2}})+(a,b)(ax_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big)$
$-2\Big((a,b)(b_{0}x_{0})(ay_{\frac{1}{2}})+(b_{0}x_{0})(b_{0}y_{\frac{1}{2}})+(b_{0}x_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big)$
$-2\Big((a,b)(b_{\frac{1}{2}}x_{0})(ay_{\frac{1}{2}})+(b_{\frac{1}{2}}x_{0})(b_{0}y_{\frac{1}{2}})+(b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big)$
$\iff$
$\textstyle{\frac{1}{2}(a,b)(1-(a,b))y_{\frac{1}{2}}x_{0}+(y_{\frac{1}{2}}x_{0})b_{0}+b_{\frac{1}{2}}(y_{\frac{1}{2}}x_{0})-2\Big(\frac{1}{2}(a,b)(y_{\frac{1}{2}}x_{0})b_{0}+\frac{1}{2}(a,b)b_{\frac{1}{2}}(y_{\frac{1}{2}}x_{0})\Big)}$
$\textstyle{-2\Big(\frac{1}{2}(a,b)(y_{\frac{1}{2}}b_{0})x_{0}+((y_{\frac{1}{2}}b_{0})x_{0})b_{0}+b_{\frac{1}{2}}((y_{\frac{1}{2}}b_{0})x_{0})}$
$+((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{0})b_{0}+b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{0})\Big)$
$y_{\frac{1}{2}}(x_{0}b_{0})+y_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})$
$\textstyle{-2\Big(\frac{1}{2}(a,b)y_{\frac{1}{2}}(x_{0}b_{0})+(y_{\frac{1}{2}}b_{0})(x_{0}b_{0})+(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{0}b_{0})\Big)}$
$\textstyle{-2\Big(\frac{1}{2}(a,b)y_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})+(b_{\frac{1}{2}}x_{0})(y_{\frac{1}{2}}b_{0})+(b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big)}$
$\iff$
$\textstyle{\frac{1}{2}(a,b)(1-(a,b))y_{\frac{1}{2}}x_{0}+(1-(a,b))(y_{\frac{1}{2}}x_{0})b_{0}+(1-(a,b))b_{\frac{1}{2}}(y_{\frac{1}{2}}x_{0})}$
$\textstyle{-2\Big(\frac{1}{2}(a,b)(y_{\frac{1}{2}}b_{0})x_{0}+((y_{\frac{1}{2}}b_{0})x_{0})b_{0}+b_{\frac{1}{2}}((y_{\frac{1}{2}}b_{0})x_{0})}$
$+((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{0})b_{0}+b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{0})\Big)$
$(1-(a,b))y_{\frac{1}{2}}(x_{0}b_{0})+(1-(a,b))y_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})$
$-2\Big((y_{\frac{1}{2}}b_{0})(x_{0}b_{0})+(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{0}b_{0})\Big)$
$-2\Big((b_{\frac{1}{2}}x_{0})(y_{\frac{1}{2}}b_{0})+(b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big).$
(2) Comparing the $\frac{1}{2}$-components we get,
$\textstyle{\frac{1}{2}(a,b)(1-(a,b))y_{\frac{1}{2}}x_{0}+(1-(a,b))(y_{\frac{1}{2}}x_{0})b_{0}}$
$-(a,b)(y_{\frac{1}{2}}b_{0})x_{0}-2((y_{\frac{1}{2}}b_{0})x_{0})b_{0}-2b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{0})\quad=$
$(1-(a,b))y_{\frac{1}{2}}(x_{0}b_{0})-2(y_{\frac{1}{2}}b_{0})(b_{0}x_{0})-2(b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}}).$
(3) This follows from comparing the $1$ plus 0 components of both sides of (2). ∎
Lemma 4.5.
Let $x_{\frac{1}{2}},y_{\frac{1}{2}}\in A_{\frac{1}{2}}(a),$ then
  1. (1)
    $\textstyle{\frac{1}{2}(a,b)(1-(a,b))(x_{\frac{1}{2}},y_{\frac{1}{2}})a+(1-(a,b))(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}+(1-(a,b))b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})}$
    $-(a,b)(x_{\frac{1}{2}},y_{\frac{1}{2}}b_{0})a-2(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}-2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))$
    $-2((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}-2b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})$
    $=\textstyle{\frac{1}{2}(a,b)(1-(a,b))x_{\frac{1}{2}}y_{\frac{1}{2}}+(1-(a,b))y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})+(1-(a,b))(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}-(a,b)x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})}$
    $-2(x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})-2(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})$
    $-2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})-2(b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{\frac{1}{2}})$
  2. (2)
    Comparing the $\frac{1}{2}$-component in (1) we get
    $(1-(a,b))b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})+2(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})+2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})$
    $=(1-(a,b))(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))+2((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}.$
  3. (3)
    Interchanging $x_{\frac{1}{2}}$ and $y_{\frac{1}{2}}$ in (2) we get:
    1. (a)
      by subtracting:
      $(1-(a,b))(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))+2((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}\quad=$
      $(1-(a,b))(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}+2b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))+2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}.$
    2. (b)
      By adding:
      $2(1-(a,b))b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})+4(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})+4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})\quad=$
      $(1-(a,b))\Big[(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]+2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))+2b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))$
      $+2((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}$
      $\iff$
      $((a,b)-1)\Big[(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]+4(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})+4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})\quad=$
      $2((a,b)-1)b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})+2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))+2b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))$
      $+2((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}$
  4. (4)
    Comparing the 0-components in (1) we get
    $\textstyle{(1-(a,b))(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}-\frac{1}{2}(a,b)(1-(a,b))\Big(x_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}}$
    $+2\Big((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}+2\Big((b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big)_{0}$
    $=(1-(a,b))\Big(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}-(a,b)\Big(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})\Big)_{0}$
    $+2((x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})\Big)_{0}$
    $\iff$
    $(a,b)(1-(a,b))\Big(x_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}-4\Big((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}-4\Big((b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big)_{0}$
    $=-2(1-(a,b))\Big(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}+2(a,b)\Big(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})\Big)_{0}$
    $-4((x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}-4\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})\Big)_{0}+2(1-(a,b))(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}$
  5. (5)
    Interchanging $x_{\frac{1}{2}},y_{\frac{1}{2}}$ in (4) and subtracting we get,
    $\Big(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}+2(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})\Big)_{0}$
    $=\Big(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})\Big)_{0}+2(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))b_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})\Big)_{0}$
  6. (6)
    Interchanging $x_{\frac{1}{2}},y_{\frac{1}{2}}$ in (4) and adding we get
    $2(1-(a,b))(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}-(a,b)(1-(a,b))\Big(x_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}$
    $+4\Big((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}+4\Big((b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big)_{0}$
    $=(1-2(a,b))\Big(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}+(1-2(a,b))\Big(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})\Big)_{0}$
    $+2((x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}+2((y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))b_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})\Big)_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})\Big)_{0}$
    $\iff$
    $(a,b)((a,b)-1)\Big(x_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}+(2(a,b)-1)\Big[\Big(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}+\Big(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})\Big)_{0}\Big]$
    $+4\Big((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}+4\Big((b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big)_{0}$
    $=-2(1-(a,b))(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}+$
    $+2((x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}+2((y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))b_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})\Big)_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})\Big)_{0}$
  7. (7)
    If $(a,b)\neq 1,$ then $((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}=\\ \frac{1}{2}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]+\frac{1}{4}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-\frac{1}{4}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
  8. (8)
    If $(a,b)\neq 1,$ then $b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})+(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})\\ =\frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}+\frac{1}{8}(1-(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}\\ +\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}$
Proof.
(1) Replacing $x=x_{\frac{1}{2}},y=y_{\frac{1}{2}}$ in Lemma 4.2(2) we get
$(a,b)(1-(a,b))a(x_{\frac{1}{2}}y_{\frac{1}{2}})+b_{0}(x_{\frac{1}{2}}y_{\frac{1}{2}})+b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})-2\Big((a,b)b_{0}((ay_{\frac{1}{2}})x_{\frac{1}{2}})+(a,b)b_{\frac{1}{2}}((ay_{\frac{1}{2}})x_{\frac{1}{2}})\Big)$
$-2\Big((a,b)a((b_{0}y_{\frac{1}{2}})x_{\frac{1}{2}})+b_{0}((b_{0}y_{\frac{1}{2}})x_{\frac{1}{2}})+b_{\frac{1}{2}}((b_{0}y_{\frac{1}{2}})x_{\frac{1}{2}})\Big)$
$-2\Big((a,b)a((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})+b_{0}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})+b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})\Big)$
$=(a,b)(1-(a,b))(ax_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{0}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}-2\Big((a,b)(ax_{\frac{1}{2}})(b_{0}y_{\frac{1}{2}})+(a,b)(ax_{\frac{1}{2}})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big)$
$-2\Big((a,b)(b_{0}x_{\frac{1}{2}})(ay_{\frac{1}{2}})+(b_{0}x_{\frac{1}{2}})(b_{0}y_{\frac{1}{2}})+(b_{0}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big)$
$-2\Big((a,b)(b_{\frac{1}{2}}x_{\frac{1}{2}})(ay_{\frac{1}{2}})+(b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{0}y_{\frac{1}{2}})+(b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{\frac{1}{2}})$
$\iff$
$\overbrace{\textstyle{\frac{1}{2}(a,b)(1-(a,b))(x_{\frac{1}{2}},y_{\frac{1}{2}})a}}^{(1)}+\overbrace{(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}}^{(2)}+\overbrace{b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})}^{(3)}-2\Big(\overbrace{\textstyle{\frac{1}{2}(a,b)(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}}}^{(2)}+\overbrace{\textstyle{\frac{1}{2}(a,b)b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})}}^{(3)}\Big)$
$-2\Big(\overbrace{\textstyle{\frac{1}{2}(a,b)(x_{\frac{1}{2}},y_{\frac{1}{2}}b_{0})a}}^{(4)}+\overbrace{(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}}^{(5)}+\overbrace{b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))}^{(6)}$
$-2\Big(\overbrace{\textstyle{\frac{1}{2}(a,b)(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}}}^{(7)}+\overbrace{((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}}^{(8)}+\overbrace{b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})}^{(9)}\Big)$
$=\overbrace{\textstyle{\frac{1}{2}(a,b)(1-(a,b))x_{\frac{1}{2}}y_{\frac{1}{2}}}}^{(10)}+\overbrace{y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})}^{(11)}+\overbrace{(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}}^{(12)}-2\Big(\overbrace{\textstyle{\frac{1}{2}(a,b)x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})}}^{(13)}+\overbrace{\textstyle{\frac{1}{2}(a,b)x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{\frac{1}{2}})}}^{(7)}\Big)$
$-2\Big(\overbrace{\textstyle{\frac{1}{2}(a,b)y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})}}^{(14)}+(x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})+(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})\Big)$
$-2\Big(\overbrace{\textstyle{\frac{1}{2}(a,b)(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}}}^{(12)}+(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})+(b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big)$
$\iff$
$\overbrace{\textstyle{\frac{1}{2}(a,b)(1-(a,b))(x_{\frac{1}{2}},y_{\frac{1}{2}})a}}^{(1)}+\overbrace{(1-(a,b))(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}}^{(2)}+\overbrace{(1-(a,b))b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})}^{(3)}$
$\overbrace{-(a,b)(x_{\frac{1}{2}},y_{\frac{1}{2}}b_{0})a}^{(4)}\overbrace{-2(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}}^{(5)}\overbrace{-2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))}^{(6)}$
$-2\overbrace{((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}}^{(8)}-2\overbrace{b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})}^{(9)}$
$=\overbrace{\textstyle{\frac{1}{2}(a,b)(1-(a,b))x_{\frac{1}{2}}y_{\frac{1}{2}}}}^{(10)}+\overbrace{y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})}^{(11)}+\overbrace{(1-(a,b))(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}}^{(12)}-\overbrace{(a,b)x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})}^{(13)}$
$\overbrace{-(a,b)y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})}^{(14)}-2(x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})-2(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})$
$-2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})-2(b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{\frac{1}{2}}).$
(2) Comparing the $\frac{1}{2}$-component in (1) we get
$(1-(a,b))b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})-2b_{\frac{1}{2}}((y_{\frac{1}{2}}b_{0})x_{\frac{1}{2}})-2((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}$
$=\quad(1-(a,b))(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}-2(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})-2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0}).$
(3) This is obvious.
(4) Comparing the 0-components in (1) we get
$(1-(a,b))(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}-2(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}-2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})\Big)_{0}$
$=\textstyle{\frac{1}{2}(a,b)(1-(a,b))\Big(x_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}+(1-(a,b))\Big(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}-(a,b)\Big(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})\Big)_{0}}$
$-2\Big((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}-2\Big((b_{\frac{1}{2}}x_{\frac{1}{2}})(b_{\frac{1}{2}}y_{\frac{1}{2}})\Big)_{0}.$
(5) Interchanging $x_{\frac{1}{2}},y_{\frac{1}{2}}$ in (4) we get,
$(1-(a,b))\Big(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}-(a,b)\Big(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})\Big)_{0}+2(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})\Big)_{0}$
$=(1-(a,b))\Big(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})\Big)_{0}-(a,b)\Big(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}+2(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))b_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})\Big)_{0}$
$\iff$
$\Big(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}+2(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})\Big)_{0}$
$=\Big(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})\Big)_{0}+2(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))b_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})\Big)_{0}.$
(6) This is obvious.
(7) We multiply (2) by $b_{0}$ to get
$(1-(a,b))(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}}))b_{0}+2((b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0}))b_{0}+2((b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0}))b_{0}$
$=(1-(a,b))\Big[((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}\Big]+2(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})))b_{0}.$
Using Lemma 4.12(2), and Lemma 4.13(5) we get
$\textstyle{(1-(a,b))(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}}))b_{0}+\frac{1}{2}(1-(a,b))\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]}$
$=\textstyle{(1-(a,b))\Big[((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}\Big]+\frac{1}{2}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
So if $(a,b)\neq 1,$ then, by Lemma 4.12(1),
$\textstyle{\frac{1}{4}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{4}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{2}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]}$
$\textstyle{=((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
(8) We replace $x_{\frac{1}{2}}$ with $x_{\frac{1}{2}}b_{0}$ in (2). We use Lemma 4.12(10) and Lemma 4.13((5)&(7)), to get
$(1-(a,b))b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})+2(b_{\frac{1}{2}}y_{\frac{1}{2}})((x_{\frac{1}{2}}b_{0})b_{0})+2(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))(y_{\frac{1}{2}}b_{0})$
$=(1-(a,b))(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))y_{\frac{1}{2}}+2b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0}))+2((b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0}))b_{0}$
$\iff$
$\textstyle{(1-(a,b))b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))+(1-(a,b))(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})}\textstyle{+\frac{1}{2}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
$=\textstyle{\frac{1}{8}(1-(a,b)^{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+\frac{1}{4}(1-(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
$\textstyle{+\frac{1}{2}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{2}(1-(a,b))(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}}$
$\text{since $(a,b)\neq 1$}\implies$
$b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))+(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})$
$=\textstyle{\frac{1}{8}(1-(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}-\frac{1}{4}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
$\textstyle{\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}.}$
Hence by Lemma 4.13(7), the claim holds. ∎
Lemma 4.6.
Let $b\in A$ be an axis of Jordan type half. Then
  1. (1)
    4(bx)(by)(bx)y(by)x(b,y)bx(b,x)by(b,xy)b+b(xy)=0.4(bx)(by)-(bx)y-(by)x-(b,y)bx-(b,x)by-(b,xy)b+b(xy)=0.
  2. (2)
    $4(bx)(by)=4(((a,b)a+b_{0}+b_{\frac{1}{2}})x)((a,b)a+b_{0}+b_{\frac{1}{2}})y)$
    $=4(a,b)^{2}(ax)(ay)+4(a,b)(ax)(yb_{0})+4(a,b)(ay)(xb_{0})+4(a,b)(ax)(yb_{\frac{1}{2}})$
    $+4(a,b)(ay)(xb_{\frac{1}{2}})+4(xb_{0})(yb_{0})+4(xb_{\frac{1}{2}})(yb_{\frac{1}{2}})+4(xb_{0})(yb_{\frac{1}{2}})+4(yb_{0})(xb_{\frac{1}{2}})$
  3. (3)
    $(bx)y=-(a,b)(ax)y-(xb_{0})y-(xb_{\frac{1}{2}})y$ and $(by)x=-(a,b)(ay)x-(yb_{0})x-(yb_{\frac{1}{2}})x.$
  4. (4)
    $(b,y)bx=-(b,y)(a,b)ax-(b,y)xb_{0}-(b,y)xb_{\frac{1}{2}}$ and $(b,x)by=-(b,x)(a,b)ay-(b,x)yb_{0}-(b,x)yb_{\frac{1}{2}}.$
  5. (5)
    $(b,xy)b+b(xy)=-(b,xy)(a,b)a-(b,xy)b_{0}-(b,xy)b_{\frac{1}{2}}+(a,b)a(xy)+(xy)b_{0}+(xy)b_{\frac{1}{2}}.$
Proof.
(1) follows from Lemma 2.3, since $P_{b}(x,y)=0,$ for all $x,y\in A.$ Then (2)–(5) are obvious.
Remark 4.7.
In Lemmas 4.8, 4.10 and 4.13 ahead we replace $x,y$ by appropriate elements, we sum up parts (2),(3),(4) and (5) of Lemma 4.6 and equal the sum to zero.
Lemma 4.8.
Putting $x=x_{0},y=a$ in Lemma 4.6 gives
  1. (1)
    $2b_{\frac{1}{2}}(x_{0}b_{0})=(1-(a,b))b_{\frac{1}{2}}x_{0}+\frac{1}{2}(b,x_{0})b_{\frac{1}{2}}.$
  2. (2)
    $2b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})=(a,b)(b,x_{0})a+(a,b)x_{0}b_{0}.$
  3. (3)
    $2(x_{0}b_{0})b_{0}=(1-(a,b))x_{0}b_{0}+(b,x_{0})b_{0}.$
  4. (4)
    $(b_{\frac{1}{2}}x_{0})b_{0}=\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}.$
Proof.
(1)&(2): By Lemma 4.6(1),
$0=$
$4(a,b)^{2}(ax_{0})(aa)+4(a,b)(ax_{0})(ab_{0})+4(a,b)(aa)(x_{0}b_{0})+4(a,b)(ax_{0})(ab_{\frac{1}{2}})$
$+4(a,b)(aa)(x_{0}b_{\frac{1}{2}})+4(x_{0}b_{0})(ab_{0})+4(x_{0}b_{\frac{1}{2}})(ab_{\frac{1}{2}})+4(x_{0}b_{0})(ab_{\frac{1}{2}})+4(ab_{0})(x_{0}b_{\frac{1}{2}})$
$-(a,b)(ax_{0})a-(x_{0}b_{0})a-(x_{0}b_{\frac{1}{2}})a-(a,b)(aa)x_{0}-(ab_{0})x_{0}-(ab_{\frac{1}{2}})x_{0}$
$-(b,a)(a,b)ax_{0}-(b,a)x_{0}b_{0}-(b,a)x_{0}b_{\frac{1}{2}}-(b,x_{0})(a,b)aa-(b,x_{0})ab_{0}-(b,x_{0})ab_{\frac{1}{2}}$
$-(b,x_{0}a)b+b(x_{0}a)$
$=\textstyle{2(a,b)b_{\frac{1}{2}}x_{0}+2b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})+2b_{\frac{1}{2}}(x_{0}b_{0})-\frac{1}{2}b_{\frac{1}{2}}x_{0}-\frac{1}{2}b_{\frac{1}{2}}x_{0}}$
$\textstyle{-(a,b)x_{0}b_{0}-(a,b)b_{\frac{1}{2}}x_{0}-(a,b)(b,x_{0})a-\frac{1}{2}(b,x_{0})b_{\frac{1}{2}}.}$
Hence
$\textstyle{((a,b)-1)b_{\frac{1}{2}}x_{0}+2b_{\frac{1}{2}}(x_{0}b_{0})-\frac{1}{2}(b,x_{0})b_{\frac{1}{2}}=0,}$
and
$2b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})-(a,b)(b,x_{0})a-(a,b)x_{0}b_{0}=0$
(3) Using (2) and Lemma 4.1(4),
$(a,b)x_{0}b_{0}=x_{0}b_{0}+(b,x_{0})b_{0}-2(x_{0}b_{0})b_{0}\ \iff$
$2(x_{0}b_{0})b_{0}=(1-(a,b))x_{0}b_{0}+(b,x_{0})b_{0}.$
(4) By Lemma 4.2(3), $(1-(a,b))b_{\frac{1}{2}}x_{0}-2b_{\frac{1}{2}}(x_{0}b_{0})+2(b_{\frac{1}{2}}x_{0})b_{0}=0,$ hence, by (1), $\frac{1}{2}(b,x_{0})b_{\frac{1}{2}}+2(b_{\frac{1}{2}}x_{0})b_{0}=0.$ This shows (4). ∎
We summarize the results of Lemma 4.8 and part of Lemma 4.2 in the following theorem.
Theorem 4.9.
Let $x_{0}\in A_{0}(a)$ and $x_{\frac{1}{2}}\in A_{\frac{1}{2}}(a),$ then
  1. (1)
    $(x_{\frac{1}{2}}b_{0})b_{0}=\frac{1}{2}(1-(a,b))x_{\frac{1}{2}}b_{0}.$
  2. (2)
    $(x_{0}b_{0})b_{0}=\frac{1}{2}(1-(a,b))x_{0}b_{0}+\frac{1}{2}(b,x_{0})b_{0}.$
  3. (3)
    $(x_{0}b_{\frac{1}{2}})b_{\frac{1}{2}}=\frac{1}{2}(a,b)(b,x_{0})a+\frac{1}{2}(a,b)x_{0}b_{0}.$
  4. (4)
    1. (a)
      $\Big(x_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}b_{\frac{1}{2}}=\frac{1}{4}(a,b)(1-(a,b))x_{\frac{1}{2}}-\frac{1}{2}(a,b)x_{\frac{1}{2}}b_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}.$
    2. (b)
      $(x_{\frac{1}{2}}b_{\frac{1}{2}})b_{\frac{1}{2}}=\frac{1}{4}(a,b)(1-(a,b))x_{\frac{1}{2}}-\frac{1}{2}(a,b)x_{\frac{1}{2}}b_{0}+\frac{1}{2}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}.$
  5. (5)
    $(x_{\frac{1}{2}}b_{0})b_{\frac{1}{2}}=\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}})a+\frac{1}{4}(b,x_{\frac{1}{2}})b_{0}.$
  6. (6)
    $(x_{\frac{1}{2}}b_{\frac{1}{2}})b_{0}=\frac{1}{2}(1-(a,b))\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})b_{0}.$
  7. (7)
    $(x_{0}b_{0})b_{\frac{1}{2}}=\frac{1}{2}(1-(a,b))b_{\frac{1}{2}}x_{0}+\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}.$
  8. (8)
    $(x_{0}b_{\frac{1}{2}})b_{0}=\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}.\quad$
Proof.
(1) is Lemma 4.2(4b). (2), (3), (7) and (8) come from Lemma 4.8. (4) is Lemma 4.2(4)(c)&(d)), (5) is Lemma 4.2(5c) and (6) is Lemma 4.2(5d). ∎
Lemma 4.10.
, Putting $x=x_{0}$ and $y=y_{0}$ in Lemma 4.6 we get
  1. (1)
    $0=$
    $4(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})+4(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})-(b_{\frac{1}{2}}x_{0})y_{0}-(b_{\frac{1}{2}}y_{0})x_{0}$
    $-(b,y_{0})b_{\frac{1}{2}}x_{0}-(b,x_{0})b_{\frac{1}{2}}y_{0}-(b,x_{0}y_{0})b_{\frac{1}{2}}+b_{\frac{1}{2}}(x_{0}y_{0}).$
  2. (2)
    Hence
    $b_{\frac{1}{2}}(x_{0}y_{0})-(b_{\frac{1}{2}}x_{0})y_{0}-(b_{\frac{1}{2}}y_{0})x_{0}$
    $=-4(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})-4(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})+(b,y_{0})b_{\frac{1}{2}}x_{0}+(b,x_{0})b_{\frac{1}{2}}y_{0}+(b,x_{0}y_{0})b_{\frac{1}{2}}.$
  3. (3)
    $((b_{\frac{1}{2}}y_{0})x_{0})b_{0}+((b_{\frac{1}{2}}x_{0})y_{0})b_{0}=\frac{1}{4}(b,x_{0}y_{0})b_{\frac{1}{2}}.$
  4. (4)
    $4(b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{0})+4(x_{0}b_{0})(y_{0}b_{0})-y_{0}(x_{0}b_{0})-x_{0}(y_{0}b_{0})+(x_{0}y_{0})b_{0}$
    $=(b,y_{0})x_{0}b_{0}+(b,x_{0})y_{0}b_{0}+(b,x_{0}y_{0})b_{0}+(b,x_{0}y_{0})(a,b)a.$
Proof.
(1)&(4) By Lemma 4.6,
$0=$
$4(a,b)^{2}(ax_{0})(ay_{0})+4(a,b)(ax_{0})(y_{0}b_{0})+4(a,b)(ay_{0})(x_{0}b_{0})+4(a,b)(ax_{0})(y_{0}b_{\frac{1}{2}})$
$+4(a,b)(ay_{0})(x_{0}b_{\frac{1}{2}})+4(x_{0}b_{0})(y_{0}b_{0})+4(x_{0}b_{\frac{1}{2}})(y_{0}b_{\frac{1}{2}})+4(x_{0}b_{0})(y_{0}b_{\frac{1}{2}})+4(y_{0}b_{0})(x_{0}b_{\frac{1}{2}})$
$-(a,b)(ax_{0})y_{0}-(x_{0}b_{0})y_{0}-(x_{0}b_{\frac{1}{2}})y_{0}-(a,b)(ay_{0})x_{0}-(y_{0}b_{0})x_{0}-(y_{0}b_{\frac{1}{2}})x_{0}$
$-(b,y_{0})(a,b)ax_{0}-(b,y_{0})(x_{0}b_{0})-(b,y_{0})x_{0}b_{\frac{1}{2}}-(b,x_{0})(a,b)ay_{0}-(b,x_{0})(y_{0}b_{0})-(b,x_{0})y_{0}b_{\frac{1}{2}}$
$-(b,x_{0}y_{0})(a,b)a-(b,x_{0}y_{0})b_{0}-(b,x_{0}y_{0})b_{\frac{1}{2}}+(a,b)a(x_{0}y_{0})+(x_{0}y_{0})b_{0}+(x_{0}y_{0})b_{\frac{1}{2}}$
$4(x_{0}b_{0})(y_{0}b_{0})+4(x_{0}b_{\frac{1}{2}})(y_{0}b_{\frac{1}{2}})+4(x_{0}b_{0})(y_{0}b_{\frac{1}{2}})+4(y_{0}b_{0})(x_{0}b_{\frac{1}{2}})$
$-(x_{0}b_{0})y_{0}-(x_{0}b_{\frac{1}{2}})y_{0}-(y_{0}b_{0})x_{0}-(y_{0}b_{\frac{1}{2}})x_{0}$
$-(b,y_{0})(x_{0}b_{0})-(b,y_{0})x_{0}b_{\frac{1}{2}}-(b,x_{0})(y_{0}b_{0})-(b,x_{0})y_{0}b_{\frac{1}{2}}$
$-(b,x_{0}y_{0})(a,b)a-(b,x_{0}y_{0})b_{0}-(b,x_{0}y_{0})b_{\frac{1}{2}}+(x_{0}y_{0})b_{0}+(x_{0}y_{0})b_{\frac{1}{2}}$
Hence
$4(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})+4(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})-(b_{\frac{1}{2}}x_{0})y_{0}-(b_{\frac{1}{2}}y_{0})x_{0}$
$-(b,y_{0})b_{\frac{1}{2}}x_{0}-(b,x_{0})b_{\frac{1}{2}}y_{0}-(b,x_{0}y_{0})b_{\frac{1}{2}}+b_{\frac{1}{2}}(x_{0}y_{0})=0.$
This shows (1) and similarly (4) holds.
(2) This immediate from (1).
(3) By Lemma 4.2(8b) we have,
$2b_{\frac{1}{2}}(x_{0}y_{0})+4(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})+4(b_{\frac{1}{2}}x_{0})(b_{0}y_{0})$
$=(1+(a,b))(b_{\frac{1}{2}}x_{0})y_{0}+(1+(a,b))(b_{\frac{1}{2}}y_{0})x_{0}+2b_{\frac{1}{2}}((y_{0}b_{0})x_{0})+2b_{\frac{1}{2}}((x_{0}b_{0})y_{0})$
$+2((b_{\frac{1}{2}}y_{0})x_{0})b_{0}+2((b_{\frac{1}{2}}x_{0})y_{0})b_{0}.$
Hence, by (2),
$b_{\frac{1}{2}}(x_{0}y_{0})-4(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})-4(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})+(b,y_{0})b_{\frac{1}{2}}x_{0}+(b,x_{0})b_{\frac{1}{2}}y_{0}+(b,x_{0}y_{0})b_{\frac{1}{2}}$
$+4(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})+4(b_{\frac{1}{2}}x_{0})(b_{0}y_{0})$
$=(a,b)(b_{\frac{1}{2}}x_{0})y_{0}+(a,b)(b_{\frac{1}{2}}y_{0})x_{0}+2b_{\frac{1}{2}}((y_{0}b_{0})x_{0})+2b_{\frac{1}{2}}((x_{0}b_{0})y_{0})$
$+2((b_{\frac{1}{2}}y_{0})x_{0})b_{0}+2((b_{\frac{1}{2}}x_{0})y_{0})b_{0}$
$\iff$
$b_{\frac{1}{2}}(x_{0}y_{0})+(b,y_{0})b_{\frac{1}{2}}x_{0}+(b,x_{0})b_{\frac{1}{2}}y_{0}+(b,x_{0}y_{0})b_{\frac{1}{2}}$
$=(a,b)(b_{\frac{1}{2}}x_{0})y_{0}+(a,b)(b_{\frac{1}{2}}y_{0})x_{0}+2b_{\frac{1}{2}}((y_{0}b_{0})x_{0})+2b_{\frac{1}{2}}((x_{0}b_{0})y_{0})$
$+2((b_{\frac{1}{2}}y_{0})x_{0})b_{0}+2((b_{\frac{1}{2}}x_{0})y_{0})b_{0}.$
Multiplying by $b_{0}$ using $(z_{\frac{1}{2}}b_{0})b_{0}=\frac{1}{2}(1-(a,b))z_{\frac{1}{2}}b_{0},\ \forall z_{\frac{1}{2}}\in A_{\frac{1}{2}}(a),$ and $(b_{\frac{1}{2}}z_{0})b_{0}=\frac{1}{4}(b,z_{0})b_{0},\ \forall z_{0}\in A_{0}(a),$ we get
$\textstyle{\frac{1}{4}(b,x_{0}y_{0})b_{\frac{1}{2}}+\frac{1}{2}(b,y_{0})(b,x_{0})b_{\frac{1}{2}}+\frac{1}{2}(1-(a,b))(b,x_{0}y_{0})b_{\frac{1}{2}}}$
$=(a,b)(((b_{\frac{1}{2}}x_{0})y_{0}))b_{0}+(a,b)((b_{\frac{1}{2}}y_{0})x_{0})b_{0}+(b,(y_{0}b_{0})x_{0})b_{\frac{1}{2}}$
$+(1-(a,b))((b_{\frac{1}{2}}y_{0})x_{0})b_{0}+(1-(a,b))(((b_{\frac{1}{2}}x_{0})y_{0}))b_{0}.$
Now we use $(y_{0}b_{0})b_{0}=\frac{1}{2}(1-(a,b))y_{0}b_{0}+\frac{1}{2}(b,y_{0})b_{0}$ to see that
$\textstyle{(b,(y_{0}b_{0})x_{0})=((y_{0}b_{0})b_{0},x_{0})=\frac{1}{2}(1-(a,b))(y_{0}b_{0},x_{0})+\frac{1}{2}(b,y_{0})(b_{0},x_{0})}$
$=\textstyle{\frac{1}{2}(1-(a,b))(b_{0},x_{0}y_{0})+\frac{1}{2}(b,x_{0})(b,y_{0}),}$
so after cancellation we are left with $\frac{1}{4}(b,x_{0}y_{0})b_{\frac{1}{2}}=((b_{\frac{1}{2}}y_{0})x_{0})b_{0}+((b_{\frac{1}{2}}x_{0})y_{0})b_{0}.$
Lemma 4.11.
Putting $x=x_{\frac{1}{2}}$ and $y=y_{0}$ in Lemma 4.6 we get
  1. (1)
    $0=$
    $(2(a,b)-1)x_{\frac{1}{2}}(y_{0}b_{0})+4(x_{\frac{1}{2}}b_{0})(y_{0}b_{0})+4(x_{\frac{1}{2}}b_{\frac{1}{2}})(b_{\frac{1}{2}}y_{0})-(x_{\frac{1}{2}}b_{0})y_{0}$
    $\textstyle{-\frac{1}{2}(b,y_{0})(a,b)x_{\frac{1}{2}}-(b,y_{0})x_{\frac{1}{2}}b_{0}-(b,x_{\frac{1}{2}})b_{\frac{1}{2}}y_{0}-(b,x_{\frac{1}{2}}y_{0})b_{\frac{1}{2}}+(x_{\frac{1}{2}}y_{0})b_{0}.}$
  2. (2)
    $0=$
    $b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})+(2(a,b)-1)(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+4(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})$
    $+4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-(b,y_{0})x_{\frac{1}{2}}b_{\frac{1}{2}}-(b,x_{\frac{1}{2}})y_{0}b_{0}-(b,x_{\frac{1}{2}}y_{0})(a,b)a-(b,x_{\frac{1}{2}}y_{0})b_{0}.$
Proof.
By Lemma 4.6, replacing $x$ with $x_{\frac{1}{2}}$ and $y$ with $y_{0},$
$0=$
$4(a,b)^{2}(ax_{\frac{1}{2}})(ay_{0})+4(a,b)(ax_{\frac{1}{2}})(y_{0}b_{0})+4(a,b)(ay_{0})(x_{\frac{1}{2}}b_{0})+4(a,b)(ax_{\frac{1}{2}})(y_{0}b_{\frac{1}{2}})$
$+4(a,b)(ay_{0})(x_{\frac{1}{2}}b_{\frac{1}{2}})+4(x_{\frac{1}{2}}b_{0})(y_{0}b_{0})+4(x_{\frac{1}{2}}b_{\frac{1}{2}})(y_{0}b_{\frac{1}{2}})+4(x_{\frac{1}{2}}b_{0})(y_{0}b_{\frac{1}{2}})+4(y_{0}b_{0})(x_{\frac{1}{2}}b_{\frac{1}{2}})$
$-(a,b)(ax_{\frac{1}{2}})y_{0}-(x_{\frac{1}{2}}b_{0})y_{0}-(x_{\frac{1}{2}}b_{\frac{1}{2}})y_{0}-(a,b)(ay_{0})x_{\frac{1}{2}}-(y_{0}b_{0})x_{\frac{1}{2}}-(y_{0}b_{\frac{1}{2}})x_{\frac{1}{2}}$
$-(b,y_{0})(a,b)ax_{\frac{1}{2}}-(b,y_{0})x_{\frac{1}{2}}b_{0}-(b,y_{0})x_{\frac{1}{2}}b_{\frac{1}{2}}-(b,x_{\frac{1}{2}})(a,b)ay_{0}-(b,x_{\frac{1}{2}})y_{0}b_{0}-(b,x_{\frac{1}{2}})y_{0}b_{\frac{1}{2}}$
$-(b,x_{\frac{1}{2}}y_{0})(a,b)a-(b,x_{\frac{1}{2}}y_{0})b_{0}-(b,x_{\frac{1}{2}}y_{0})b_{\frac{1}{2}}+(a,b)a(x_{\frac{1}{2}}y_{0})+(x_{\frac{1}{2}}y_{0})b_{0}+(x_{\frac{1}{2}}y_{0})b_{\frac{1}{2}}$
$2(a,b)x_{\frac{1}{2}}(y_{0}b_{0})+2(a,b)x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0})$
$+4(x_{\frac{1}{2}}b_{0})(y_{0}b_{0})+4(x_{\frac{1}{2}}b_{\frac{1}{2}})(b_{\frac{1}{2}}y_{0})+4(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})+4(y_{0}b_{0})(x_{\frac{1}{2}}b_{\frac{1}{2}})$
$\textstyle{-\frac{1}{2}(a,b)x_{\frac{1}{2}}y_{0}-(x_{\frac{1}{2}}b_{0})y_{0}-(x_{\frac{1}{2}}b_{\frac{1}{2}})y_{0}-x_{\frac{1}{2}}(y_{0}b_{0})-(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}}$
$\textstyle{-\frac{1}{2}(b,y_{0})(a,b)x_{\frac{1}{2}}-(b,y_{0})x_{\frac{1}{2}}b_{0}-(b,y_{0})x_{\frac{1}{2}}b_{\frac{1}{2}}-(b,x_{\frac{1}{2}})y_{0}b_{0}-(b,x_{\frac{1}{2}})b_{\frac{1}{2}}y_{0}}$
$\textstyle{-(b,x_{\frac{1}{2}}y_{0})(a,b)a-(b,x_{\frac{1}{2}}y_{0})b_{0}-(b,x_{\frac{1}{2}}y_{0})b_{\frac{1}{2}}+\frac{1}{2}(a,b)x_{\frac{1}{2}}y_{0}+(x_{\frac{1}{2}}y_{0})b_{0}+(x_{\frac{1}{2}}y_{0})b_{\frac{1}{2}}.}$
Hence, since the $\frac{1}{2}$-component is $0,$
$0=$
$2(a,b)x_{\frac{1}{2}}(y_{0}b_{0})+4(x_{\frac{1}{2}}b_{0})(y_{0}b_{0})+4(x_{\frac{1}{2}}b_{\frac{1}{2}})(b_{\frac{1}{2}}y_{0})$
$\textstyle{-\frac{1}{2}(a,b)x_{\frac{1}{2}}y_{0}-(x_{\frac{1}{2}}b_{0})y_{0}-x_{\frac{1}{2}}(y_{0}b_{0})}$
$\textstyle{-\frac{1}{2}(b,y_{0})(a,b)x_{\frac{1}{2}}-(b,y_{0})x_{\frac{1}{2}}b_{0}-(b,x_{\frac{1}{2}})b_{\frac{1}{2}}y_{0}}$
$\textstyle{-(b,x_{\frac{1}{2}}y_{0})b_{\frac{1}{2}}+\frac{1}{2}(a,b)x_{\frac{1}{2}}y_{0}+(x_{\frac{1}{2}}y_{0})b_{0}.}$
And, since the $1$-component plus the 0-compoments are $0,$
$0=$
$2(a,b)x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0})+4(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})+4(y_{0}b_{0})(x_{\frac{1}{2}}b_{\frac{1}{2}})$
$-(x_{\frac{1}{2}}b_{\frac{1}{2}})y_{0}-(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}-(b,y_{0})x_{\frac{1}{2}}b_{\frac{1}{2}}-(b,x_{\frac{1}{2}})y_{0}b_{0}$
$-(b,x_{\frac{1}{2}}y_{0})(a,b)a-(b,x_{\frac{1}{2}}y_{0})b_{0}+(x_{\frac{1}{2}}y_{0})b_{\frac{1}{2}}.\qed$
Lemma 4.12.
  1. (1)
    $(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}}))b_{0}=\frac{1}{4}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{4}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
  2. (2)
    $(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}=\frac{1}{4}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
  3. (3)
    $(a,(x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})=\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}}).$
  4. (4)
    $(b_{0},(x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})=\frac{1}{2}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}}).$
  5. (5)
    $(a,(x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0}))=\frac{1}{4}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}}).$
  6. (6)
    $(b_{0},(x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0}))=\frac{1}{4}(1-(a,b))^{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}}).$
  7. (7)
    $(b,(x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})=\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}}).$
  8. (8)
    $(b,(x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0}))=\frac{1}{4}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}}).$
  9. (9)
    $(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))y_{\frac{1}{2}}=\frac{1}{8}(1-(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+\frac{1}{4}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}.$
  10. (10)
    $(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))(y_{\frac{1}{2}}b_{0})=\frac{1}{4}(1-(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}.$
Proof.
(1) By Theorem 4.9(1), $(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}}))b_{0}=\Big(\Big[(a,x_{\frac{1}{2}}y_{\frac{1}{2}})a+\Big(x_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}\Big]b_{\frac{1}{2}}\Big)b_{0}\\ \textstyle{=\frac{1}{4}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{4}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
(2) Replacing $y_{\frac{1}{2}}$ with $y_{\frac{1}{2}}b_{0}$ in (1), and using Theorem 4.9(1), we see that ${(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{0}=\frac{1}{4}(1-(a,b))(a,x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{\frac{1}{2}}+\frac{1}{4}(b_{0},x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))b_{\frac{1}{2}}}\\ =\textstyle{\frac{1}{4}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
(3) is clear and (4) follows from Theorem 4.9(1). (5) follows from (3) replacing $y_{\frac{1}{2}}$ with $y_{\frac{1}{2}}b_{0}.$ Similarly (6) follows from (4). For (7) we have $(b,(x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})=\frac{1}{2}(a,b)(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})+\frac{1}{2}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})=\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}}),$ and (8) follow from (7).
(9)&(10) $(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})))y_{\frac{1}{2}}=\Big[\frac{1}{4}(1-(a,b)(b,x_{\frac{1}{2}})a+\frac{1}{4}(b,x_{\frac{1}{2}})b_{0}\Big]y_{\frac{1}{2}}=\frac{1}{8}(1-(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+\frac{1}{4}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}.$ This shows (9), and (10) is obtained by replacing in (9) $y_{\frac{1}{2}}$ with $y_{\frac{1}{2}}b_{0}.$
Lemma 4.13.
For all $x_{\frac{1}{2}},y_{\frac{1}{2}}\in A_{\frac{1}{2}}(a),$
  1. (1)
    $0=$
    $(a,b)((a,b)-1)x_{\frac{1}{2}}y_{\frac{1}{2}}+(2(a,b)-1)x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})+(2(a,b)-1)y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})$
    $+4(x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})+4(x_{\frac{1}{2}}b_{\frac{1}{2}})(y_{\frac{1}{2}}b_{\frac{1}{2}})$
    $-(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{\frac{1}{2}}-(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{\frac{1}{2}}$
    $-(b,x_{\frac{1}{2}}y_{\frac{1}{2}})(a,b)a-(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}+(a,b)a(x_{\frac{1}{2}}y_{\frac{1}{2}})+(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}$
    $\iff$
    $(2(a,b)-1)\Big(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})\Big)_{0}+(2(a,b)-1)\Big(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}$
    $-(b,y_{\frac{1}{2}})\Big(x_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}-(b,x_{\frac{1}{2}})\Big(y_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}+(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}-(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}\qquad=$
    $(a,b)(1-(a,b))\Big(x_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}-4\Big((x_{\frac{1}{2}}b_{\frac{1}{2}})(y_{\frac{1}{2}}b_{\frac{1}{2}})\Big)_{0}-4\Big((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}.$
  2. (2)
    $0=$
    $+(2(a,b)-1)\Big[(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]+4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})+4(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})$
    $\textstyle{-\frac{1}{2}(a,b)\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}\Big]-(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}-(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
    $-(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}$
    $\iff$
    $(2(a,b)-1)\Big[(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]+4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})+4(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})\quad=$
    $\textstyle{\frac{1}{2}(a,b)\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}\Big]+(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
    $+(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
  3. (3)
    $-(a,b)\Big[(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]$
    $\textstyle{+\frac{1}{2}(a,b)\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}\Big]+(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
    $+(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}$
    $+2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))+2b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))$
    $+2((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}$
    $+(2(a,b)-1)b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})$
  4. (4)
    (4.1) $\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]-2\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}\Big]$
    $\textstyle{=\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-\frac{1}{2}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}}$
    $\textstyle{=\frac{1}{2}(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-\frac{1}{2}(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
  5. (5)
    $((b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0}))b_{0}=\frac{1}{4}(1-(a,b))(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}.$ Hence,
    $\Big[\Big(b_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}(x_{\frac{1}{2}}b_{0})\Big]b_{0}=\frac{1}{4}(1-(a,b))(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}-\frac{1}{4}(b,y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})b_{0}=\frac{1}{8}(1-(a,b))(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}$
  6. (6)
    $(1-2(a,b))b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})-(a,b)\Big[(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]$
    $\textstyle{+\frac{1}{2}(a,b)\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}\Big]}$
    $+\textstyle{\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-\frac{1}{2}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}}$
    $2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))+2b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})).$
  7. (7)
    1. (a)
      $b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0}))=\frac{1}{4}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
    2. (b)
      $((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0}))b_{0}=\frac{1}{4}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}$
  8. (8)
    If $(a,b)\neq 0,$ then
    $b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))+(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})$
    $\textstyle{=\frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}+\frac{1}{8}(1-(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}}$
    $\textstyle{+\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
Proof.
(1)&(2) By Lemma 4.6,
$0=$
$4(a,b)^{2}(ax_{\frac{1}{2}})(ay_{\frac{1}{2}})+4(a,b)(ax_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})+4(a,b)(ay_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})+4(a,b)(ax_{\frac{1}{2}})(y_{\frac{1}{2}}b_{\frac{1}{2}})$
$+4(a,b)(ay_{\frac{1}{2}})(x_{\frac{1}{2}}b_{\frac{1}{2}})+4(x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})+4(x_{\frac{1}{2}}b_{\frac{1}{2}})(y_{\frac{1}{2}}b_{\frac{1}{2}})+4(x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{\frac{1}{2}})+4(y_{\frac{1}{2}}b_{0})(x_{\frac{1}{2}}b_{\frac{1}{2}})$
$-(a,b)(ax_{\frac{1}{2}})y_{\frac{1}{2}}-(x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}}-(x_{\frac{1}{2}}b_{\frac{1}{2}})y_{\frac{1}{2}}-(a,b)(ay_{\frac{1}{2}})x_{\frac{1}{2}}-(y_{\frac{1}{2}}b_{0})x_{\frac{1}{2}}-(y_{\frac{1}{2}}b_{\frac{1}{2}})x_{\frac{1}{2}}$
$-(b,y_{\frac{1}{2}})(a,b)ax_{\frac{1}{2}}-(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}-(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{\frac{1}{2}}-(b,x_{\frac{1}{2}})(a,b)ay_{\frac{1}{2}}-(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}-(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{\frac{1}{2}}$
$-(b,x_{\frac{1}{2}}y_{\frac{1}{2}})(a,b)a-(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}-(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+(a,b)a(x_{\frac{1}{2}}y_{\frac{1}{2}})+(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}+(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}$
$(a,b)^{2}x_{\frac{1}{2}}y_{\frac{1}{2}}+2(a,b)x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})+2(a,b)y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})+2(a,b)x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{\frac{1}{2}})+2(a,b)y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{\frac{1}{2}})$
$+4(x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})+4(x_{\frac{1}{2}}b_{\frac{1}{2}})(y_{\frac{1}{2}}b_{\frac{1}{2}})+4(x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{\frac{1}{2}})+4(y_{\frac{1}{2}}b_{0})(x_{\frac{1}{2}}b_{\frac{1}{2}})$
$\textstyle{-\frac{1}{2}(a,b)x_{\frac{1}{2}}y_{\frac{1}{2}}-(x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}}-(x_{\frac{1}{2}}b_{\frac{1}{2}})y_{\frac{1}{2}}-\frac{1}{2}(a,b)y_{\frac{1}{2}}x_{\frac{1}{2}}-(y_{\frac{1}{2}}b_{0})x_{\frac{1}{2}}-(y_{\frac{1}{2}}b_{\frac{1}{2}})x_{\frac{1}{2}}}$
$\textstyle{-\frac{1}{2}(a,b)(b,y_{\frac{1}{2}})x_{\frac{1}{2}}-(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}-(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{\frac{1}{2}}-\frac{1}{2}(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}-(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}-(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{\frac{1}{2}}}$
$-(b,x_{\frac{1}{2}}y_{\frac{1}{2}})(a,b)a-(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}-(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+(a,b)a(x_{\frac{1}{2}}y_{\frac{1}{2}})+(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}+(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}$
$(a,b)((a,b)-1)x_{\frac{1}{2}}y_{\frac{1}{2}}+(2(a,b)-1)x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})+(2(a,b)-1)y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})$
$+(2(a,b)-1)x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{\frac{1}{2}})+(2(a,b)-1)y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{\frac{1}{2}})$
$+4(x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})+4(x_{\frac{1}{2}}b_{\frac{1}{2}})(y_{\frac{1}{2}}b_{\frac{1}{2}})+4(x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{\frac{1}{2}})+4(y_{\frac{1}{2}}b_{0})(x_{\frac{1}{2}}b_{\frac{1}{2}})$
$\textstyle{-\frac{1}{2}(a,b)(b,y_{\frac{1}{2}})x_{\frac{1}{2}}-(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}-(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{\frac{1}{2}}-\frac{1}{2}(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}-(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}-(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{\frac{1}{2}}}$
$-(b,x_{\frac{1}{2}}y_{\frac{1}{2}})(a,b)a-(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}-(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+(a,b)a(x_{\frac{1}{2}}y_{\frac{1}{2}})+(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}+(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
Thus the $\{0,1\}$ component is $0,$ which is (1), and the $\frac{1}{2}$-component is $0,$ which is (2).∎
(3) To prove (3) we use Lemma 4.5(3)(b) and part (2) to get:
$-(a,b)\Big[(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]$
$\textstyle{+\frac{1}{2}(a,b)\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}\Big]+(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
$+(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})$
$+2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))+2b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))$
$+2((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}$
$+2((a,b)-1)b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}}),$
and (3) follows.
(4) Multiplying (3) by $b_{0},$ using Theorem 4.9(1), we get
$-(a,b)\Big[((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}\Big]$
$\textstyle{\frac{1}{2}(a,b)\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]}$
$\textstyle{+\frac{1}{2}(1-(a,b))\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]}$
$\textstyle{+\frac{1}{2}(1-(a,b))(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}}$
$+2(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})))b_{0}+2(b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})))b_{0}+(1-(a,b))\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}\Big]$
$+(2(a,b)-1)(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}}))b_{0}$
$\iff$
$\textstyle{\frac{1}{2}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]}$
$\textstyle{+\frac{1}{2}(1-(a,b))(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}}$
$+2(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})))b_{0}+2(b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})))b_{0}$
$\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}\Big]$
$+(2(a,b)-1)(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}}))b_{0}.$
Using Lemma 4.12(1&2) we get
$\textstyle{\frac{1}{2}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]}$
$\textstyle{+\frac{1}{2}(1-(a,b))(a,b)(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{2}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}}$
$(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}$
$+\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}\Big]$
$\textstyle{(2(a,b)-1)\Big[\frac{1}{4}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{4}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}\Big]}$
$\iff$
$\textstyle{\frac{1}{2}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]-\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}\Big]}$
$\textstyle{-\frac{1}{4}(1-(a,b))(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{4}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
(5) Putting $x_{\frac{1}{2}}b_{0}$ in place of $x_{\frac{1}{2}}$ in equation (4.1) we get (using also Lemma 4.12(9)),
$\Big[(b,y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})b_{0}+(b,x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}}b_{0}\Big]-2\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0}))b_{0}+((b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))y_{\frac{1}{2}})b_{0}\Big]$
$=\textstyle{\frac{1}{2}(b_{0},(x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})b_{\frac{1}{2}}-\frac{1}{2}(1-(a,b))(a,(x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})b_{\frac{1}{2}}}$
$\iff$
$\textstyle{\frac{1}{2}(1-(a,b))\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]-2\Big[((b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0}))b_{0}+\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]}$
$=\textstyle{\frac{1}{4}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-\frac{1}{4}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}}$
$\iff$
$\textstyle{((b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0}))b_{0}=\frac{1}{4}(1-(a,b))(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}.}$
(6) This follows from (3) and (4).
(7) (a) Replacing in (6), $x_{\frac{1}{2}},y_{\frac{1}{2}}$ with $x_{\frac{1}{2}}b_{0},y_{\frac{1}{2}}b_{0},$ using Lemma 4.12 (5), (6), (8) and (10), we get
$\textstyle{(1-2(a,b))b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0}))-\frac{1}{4}(1-(a,b))(a,b)\Big[(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}+(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}\Big]}$
$\textstyle{+\frac{1}{4}(1-(a,b))(a,b)\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]}$
$\textstyle{+\frac{1}{8}(1-(a,b)^{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-\frac{1}{8}(1-(a,b))^{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{4}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}}$
$2(1-(a,b))b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})).$
(b) Replacing $x_{\frac{1}{2}},y_{\frac{1}{2}}$ by $x_{\frac{1}{2}}b_{0},y_{\frac{1}{2}}b_{0}$ in (1), and using Theorem 4.9((1)&(5)) and Lemma 4.12(8), we get,
$(2(a,b)-1)(1-(a,b))\Big((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}$
$-(b,(y_{\frac{1}{2}}b_{0}))\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}-(b,(x_{\frac{1}{2}}b_{0}))\Big(b_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})\Big)_{0}$
$+((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0}))b_{0}-(b,(x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0}))b_{0}\qquad=$
$(a,b)(1-(a,b))\Big((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}-4\Big((b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))(b_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))\Big)_{0}$
$-(1-(a,b))^{2}\Big((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}$
$\iff$
$\textstyle{((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0}))b_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})(b,y_{\frac{1}{2}})b_{0}-\frac{1}{4}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})=}$
$\textstyle{-\frac{1}{4}(b,x_{\frac{1}{2}})(b,y_{\frac{1}{2}})b_{0}.}$
(8) Replacing in (6) $x_{\frac{1}{2}}$ with $x_{\frac{1}{2}}b_{0},$ using Lemma 4.12 and part (7) we get
$\textstyle{(1-2(a,b))b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})-\frac{1}{8}(1-(a,b)(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}-\frac{1}{4}(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
$\textstyle{-(a,b)(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})+\frac{1}{2}(a,b)(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+\frac{1}{4}(1-(a,b))(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}}$
$\textstyle{+\frac{1}{4}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-\frac{1}{4}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}}$
$\textstyle{\frac{1}{2}(1-(a,b))(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+(1-(a,b))b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))}$
$\iff$
$\textstyle{-(a,b))b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})+\frac{1}{8}(1-(a,b)(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}-\frac{1}{4}(a,b)(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
$\textstyle{-(a,b)(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})+\frac{1}{2}(a,b)(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+\frac{1}{2}(a,b)(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}}$
$=0.$
5. The proof that $b_{\frac{1}{2}}(x_{0}y_{0})=(b_{\frac{1}{2}}x_{0})y_{0}+(b_{\frac{1}{2}}y_{0})x_{0}$
In this section we prove.
Theorem 5.1.
If $(a,b)\neq\frac{1}{4},$ then $b_{\frac{1}{2}}(x_{0}y_{0})=(b_{\frac{1}{2}}x_{0})y_{0}+x_{0}(b_{\frac{1}{2}}y_{0}).$
Proof.
We first show the following.
(5.1) $2(a,b)\Big((b_{\frac{1}{2}}x_{0})y_{0}+(b_{\frac{1}{2}}y_{0})x_{0}-b_{\frac{1}{2}}(x_{0}y_{0})\Big)$
$+(b,x_{0})b_{\frac{1}{2}}y_{0}+4((b_{\frac{1}{2}}x_{0})y_{0})b_{0}-4(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})=0.$
Proof of (5.1): We replace $x_{\frac{1}{2}}$ with $b_{\frac{1}{2}}x_{0}$ in Lemma 4.11(1), using $(b_{\frac{1}{2}}x_{0})b_{0}=\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}$ and $2b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})=(a,b)(b,x_{0})a+(a,b)x_{0}b_{0},$ and $b_{\frac{1}{2}}(y_{0}b_{0})=\frac{1}{2}(1-(a,b))b_{\frac{1}{2}}y_{0}+\frac{1}{4}(b,y_{0})b_{\frac{1}{2}}$ (Theorem 4.9). We get
$0=$
$(2(a,b)-1)(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})+4((b_{\frac{1}{2}}x_{0})b_{0})(y_{0}b_{0})+4((b_{\frac{1}{2}}x_{0})b_{\frac{1}{2}})(b_{\frac{1}{2}}y_{0})-((b_{\frac{1}{2}}x_{0})b_{0})y_{0}$
$\textstyle{-\frac{1}{2}(b,y_{0})(a,b)b_{\frac{1}{2}}x_{0}-(b,y_{0})(b_{\frac{1}{2}}x_{0})b_{0}-(b,b_{\frac{1}{2}}x_{0})b_{\frac{1}{2}}y_{0}-(b,(b_{\frac{1}{2}}x_{0})y_{0})b_{\frac{1}{2}}+((b_{\frac{1}{2}}x_{0})y_{0})b_{0}}$
$(2(a,b)-1)(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})+(b,x_{0})b_{\frac{1}{2}}(y_{0}b_{0})+(a,b)(b,x_{0})b_{\frac{1}{2}}y_{0}+2(a,b)(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})$
$\textstyle{-\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}y_{0}-\frac{1}{2}(b,y_{0})(a,b)b_{\frac{1}{2}}x_{0}-\frac{1}{4}(b,x_{0})(b,y_{0})b_{\frac{1}{2}}-(a,b)(b,x_{0})b_{\frac{1}{2}}y_{0}-(b,(b_{\frac{1}{2}}x_{0})y_{0})b_{\frac{1}{2}}}$
$+((b_{\frac{1}{2}}x_{0})y_{0})b_{0}$
$\textstyle{(2(a,b)-1)(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})+(b,x_{0})\Big(\frac{1}{2}(1-(a,b))b_{\frac{1}{2}}y_{0}+\frac{1}{4}(b,y_{0})b_{\frac{1}{2}}\Big)+2(a,b)(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})}$
$\textstyle{-\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}y_{0}-\frac{1}{2}(b,y_{0})(a,b)b_{\frac{1}{2}}x_{0}-\frac{1}{4}(b,x_{0})(b,y_{0})b_{\frac{1}{2}}-(b,(b_{\frac{1}{2}}x_{0})y_{0})b_{\frac{1}{2}}}$
$+((b_{\frac{1}{2}}x_{0})y_{0})b_{0}$
$\iff$
$\textstyle{2(a,b)(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})+2(a,b)(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})-\frac{1}{2}(a,b)(b,y_{0})b_{\frac{1}{2}}x_{0}-\frac{1}{2}(a,b)(b,x_{0})b_{\frac{1}{2}}y_{0}}$
$\textstyle{-\frac{1}{2}(a,b)(x_{0}b_{0},y_{0})b_{\frac{1}{2}}}$
$\textstyle{+\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}y_{0}+((b_{\frac{1}{2}}x_{0})y_{0})b_{0}-(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})=0}$
$\iff^{\prime}$
$(a,b)\Big(4(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})+4(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})-(b,y_{0})b_{\frac{1}{2}}x_{0}-(b,x_{0})b_{\frac{1}{2}}y_{0}-(b,x_{0}y_{0})b_{\frac{1}{2}}\Big)$
$\textstyle{+\frac{1}{2}(b,x_{0})b_{\frac{1}{2}}y_{0}+2((b_{\frac{1}{2}}x_{0})y_{0})b_{0}-2(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})=0.}$
Recall now that by Lemma 4.10(2),
$4(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})+4(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})-(b,y_{0})b_{\frac{1}{2}}x_{0}-(b,x_{0})b_{\frac{1}{2}}y_{0}-(b,x_{0}y_{0})b_{\frac{1}{2}}$
$=(b_{\frac{1}{2}}x_{0})y_{0}+(b_{\frac{1}{2}}y_{0})x_{0}-b_{\frac{1}{2}}(x_{0}y_{0}).$
Plugging this into the last identity gives (5.1).∎
We now complete the proof of the theorem. Interchanging $x_{0},y_{0}$ in (5.1) and adding, we get
$0=$
$4(a,b)\Big((b_{\frac{1}{2}}x_{0})y_{0}+(b_{\frac{1}{2}}y_{0})x_{0}-b_{\frac{1}{2}}(x_{0}y_{0})\Big)$
$+(b,x_{0})b_{\frac{1}{2}}y_{0}+4((b_{\frac{1}{2}}x_{0})y_{0})b_{0}-4(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})$
$+(b,y_{0})b_{\frac{1}{2}}x_{0}+4((b_{\frac{1}{2}}y_{0})x_{0})b_{0}-4(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})$
$4(a,b)\Big((b_{\frac{1}{2}}x_{0})y_{0}+(b_{\frac{1}{2}}y_{0})x_{0}-b_{\frac{1}{2}}(x_{0}y_{0})\Big)$
$+4((b_{\frac{1}{2}}x_{0})y_{0})b_{0}+4((b_{\frac{1}{2}}y_{0})x_{0})b_{0}$
$-4(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})-4(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})+(b,y_{0})b_{\frac{1}{2}}x_{0}+(b,x_{0})b_{\frac{1}{2}}y_{0}$
(using Lemma 4.10(3), and Lemma 4.10(2))
$4(a,b)\Big((b_{\frac{1}{2}}x_{0})y_{0}+(b_{\frac{1}{2}}y_{0})x_{0}-b_{\frac{1}{2}}(x_{0}y_{0})\Big)$
$+(b,x_{0}y_{0})b_{\frac{1}{2}}$
$+\Big(b_{\frac{1}{2}}x_{0})y_{0}-(b_{\frac{1}{2}}y_{0})x_{0}-b_{\frac{1}{2}}(x_{0}y_{0}\Big)-(b,x_{0}y_{0})b_{\frac{1}{2}},$
which completes the proof of the theorem.
6. The proof of the identity
$b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}=\frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}$
In this section we prove,
Theorem 6.1.
If $(a,b)\notin\{1,\frac{1}{4}\},$ then
$b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}=\\ \frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
We start with,
Lemma 6.2.
  1. (1)
    If $(a,b)\neq\frac{1}{4},$ then
    $(1-(a,b))(y_{\frac{1}{2}}x_{0})b_{0}-2((y_{\frac{1}{2}}b_{0})x_{0})b_{0}$
    $=(1-(a,b))y_{\frac{1}{2}}(x_{0}b_{0})-2(y_{\frac{1}{2}}b_{0})(b_{0}x_{0}).$
  2. (2)
    If $(a,b)\notin\{1,\frac{1}{4}\},$ then
    $((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}+(y_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}x_{\frac{1}{2}})$
    $\textstyle{=\frac{1}{2}(1-(a,b))(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}-\frac{1}{8}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+\frac{3}{4}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}.}$
  3. (3)
    If $(a,b)\notin\{1,\frac{1}{4}\},$ then
    $(y_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}x_{\frac{1}{2}})+(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}})$
    $\textstyle{=\frac{1}{2}(1-(a,b))(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+\frac{1}{2}(1-(a,b))(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}}$
    $\textstyle{-\frac{1}{8}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}-\frac{1}{8}(1-(a,b))(b,y_{\frac{1}{2}})x_{\frac{1}{2}}}$
    $\textstyle{\frac{1}{4}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]}$
    $\textstyle{+\frac{1}{4}(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-\frac{1}{4}(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
Proof.
(1) By Lemma 4.4(2) and Theorem 5.1. using also Theorem 4.9(4a), we get,
$\textstyle{\frac{1}{2}(a,b)(1-(a,b))y_{\frac{1}{2}}x_{0}+(1-(a,b))(y_{\frac{1}{2}}x_{0})b_{0}-(a,b)(y_{\frac{1}{2}}b_{0})x_{0}-2((y_{\frac{1}{2}}b_{0})x_{0})b_{0}}$
$-2b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{0})$
$=(1-(a,b))y_{\frac{1}{2}}(x_{0}b_{0})-2(y_{\frac{1}{2}}b_{0})(b_{0}x_{0})-2(b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}})$
$\iff$
$\textstyle{\frac{1}{2}(a,b)(1-(a,b))y_{\frac{1}{2}}x_{0}+(1-(a,b))(y_{\frac{1}{2}}x_{0})b_{0}-(a,b)(y_{\frac{1}{2}}b_{0})x_{0}-2((y_{\frac{1}{2}}b_{0})x_{0})b_{0}}$
$-2\Big[b_{\frac{1}{2}}\Big(b_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}\Big]x_{0}-2(b_{\frac{1}{2}}x_{0})\Big(b_{\frac{1}{2}}y_{\frac{1}{2}}\Big)_{0}$
$=(1-(a,b))y_{\frac{1}{2}}(x_{0}b_{0})-2(y_{\frac{1}{2}}b_{0})(b_{0}x_{0})-2(b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}})$
$\iff$
$\textstyle{\frac{1}{2}(a,b)(1-(a,b))y_{\frac{1}{2}}x_{0}+(1-(a,b))(y_{\frac{1}{2}}x_{0})b_{0}-(a,b)(y_{\frac{1}{2}}b_{0})x_{0}-2((y_{\frac{1}{2}}b_{0})x_{0})b_{0}}$
$\textstyle{-2\Big[\frac{1}{4}(a,b)(1-(a,b))y_{\frac{1}{2}}-\frac{1}{2}(a,b)y_{\frac{1}{2}}b_{0}+\frac{1}{4}(b,y_{\frac{1}{2}})b_{\frac{1}{2}}\Big]x_{0}+\frac{1}{2}(b,y_{\frac{1}{2}})b_{\frac{1}{2}}x_{0}}$
$=(1-(a,b))y_{\frac{1}{2}}(x_{0}b_{0})-2(y_{\frac{1}{2}}b_{0})(b_{0}x_{0})$
$\iff$
$(1-(a,b))(y_{\frac{1}{2}}x_{0})b_{0}-(a,b)(y_{\frac{1}{2}}b_{0})x_{0}-2((y_{\frac{1}{2}}b_{0})x_{0})b_{0}$
$+(a,b)x_{0}(y_{\frac{1}{2}}b_{0})$
$=(1-(a,b))y_{\frac{1}{2}}(x_{0}b_{0})-2(y_{\frac{1}{2}}b_{0})(b_{0}x_{0})$
$\iff$
$(1-(a,b))(y_{\frac{1}{2}}x_{0})b_{0}-2((y_{\frac{1}{2}}b_{0})x_{0})b_{0}$
$=(1-(a,b))y_{\frac{1}{2}}(x_{0}b_{0})-2(y_{\frac{1}{2}}b_{0})(b_{0}x_{0}).$
(2) Taking $x_{0}=\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}$ in (1), and using $\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}z=(b_{\frac{1}{2}}x_{\frac{1}{2}})z-\frac{1}{4}(b,x_{\frac{1}{2}})z,$ for $z\in\{y_{\frac{1}{2}},y_{\frac{1}{2}}b_{0}\},$ and also Lemma 4.13(5), we get
$(1-(a,b))\Big[\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}y_{\frac{1}{2}}\Big]b_{0}-2\Big[(y_{\frac{1}{2}}b_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}\Big]b_{0}$
$=\textstyle{\Big[(1-(a,b))y_{\frac{1}{2}}-2y_{\frac{1}{2}}b_{0}\Big]\Big[\frac{1}{2}(1-(a,b))\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})b_{0}\Big]}$
$\iff$
$\textstyle{(1-(a,b))\Big[(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}\Big]b_{0}-\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}-\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
$\textstyle{=\frac{1}{2}(1-(a,b))^{2}(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}-\frac{1}{8}(1-(a,b))^{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
$\textstyle{-(1-(a,b))(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})+\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
$\textstyle{-\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
$\iff$
$((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}+(y_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}x_{\frac{1}{2}})$
$\textstyle{=\frac{1}{2}(1-(a,b))(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}-\frac{1}{8}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+\frac{3}{4}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}.}$
(3) Interchanging $x_{\frac{1}{2}},y_{\frac{1}{2}}$ in (2) adding and using Equation (4.1) for $(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0},$ we get
$\textstyle{\frac{1}{2}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]-\frac{1}{4}(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+\frac{1}{4}(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}}$
$+(y_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}x_{\frac{1}{2}})+(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}})$
$\textstyle{=\frac{1}{2}(1-(a,b))\Big[(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]}$
$\textstyle{-\frac{1}{8}(1-(a,b))\Big[(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(b,y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]}$
$\textstyle{+\frac{3}{4}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}+\frac{3}{4}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}}$
$\iff$
$(y_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}x_{\frac{1}{2}})+(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{\frac{1}{2}})$
$\textstyle{=\frac{1}{2}(1-(a,b))(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+\frac{1}{2}(1-(a,b))(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}}$
$\textstyle{-\frac{1}{8}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}-\frac{1}{8}(1-(a,b))(b,y_{\frac{1}{2}})x_{\frac{1}{2}}}$
$\textstyle{+\frac{1}{4}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]}$
$\textstyle{+\frac{1}{4}(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-\frac{1}{4}(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}\qed$
Proof of Theorem 6.1 .
We replace in Lemma 4.13(2), $4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})+4(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0}),$ by the formula of Lemma 6.2(3) to get
$0=$
$(2(a,b)-1)\Big[(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]$
$+2(1-(a,b))\Big[(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]$
$\textstyle{-\frac{1}{2}(1-(a,b))\Big[(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(b,y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]}$
$(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}$
$+(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}$
$\textstyle{-\frac{1}{2}(a,b)\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}\Big]-(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}-(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
$-(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}$
$b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}$
$\textstyle{-\frac{1}{2}\Big[b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(b,y_{\frac{1}{2}})x_{\frac{1}{2}}\Big]-(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}\qed$
7. The proof of the identity
$\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}=0.$
The purpose of this section is to prove:
Theorem 7.1.
If $(a,b)\notin\{1,\frac{1}{4}\},$ then
$\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}=0.$
Proposition 7.2.
(7.1) $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}=$
$2\Big[b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big]b_{0}+$
$\textstyle{2\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-2\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}-\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}.}$
Proof.
Recall that by Lemma 4.4(3),
(7.2) $(1-(a,b))b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})-(1-(a,b))(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}+2(b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})$
$+2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}-2b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})$
$=0.$
By Lemma 4.3(3),
(7.3) $b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})+(a,b)(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+2(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})$
$+2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))$
$-2(a,b)a(x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0}))-2((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}$
$=0.$
Using $2(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0}))b_{0}-(1-(a,b))\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}=0,$ from Theorem 1.4(6), we compute (7.2) minus (7.3) to get,
$0=$
$-(a,b)\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}$
$-2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}-2\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}$
$+2\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}+2((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}$
$-(a,b)\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}$
$-2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}+2((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}$
$+2\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-2\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}$
$+2(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0}))b_{0}-(1-(a,b))\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}$
$-\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}$
$+2(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0}))b_{0}-2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}+2((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}$
$\textstyle{+2\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-2\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}-\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}.}\qed$
Lemma 7.3.
(7.4) $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}=$
$4\Big((b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}+4((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}$
$-(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}-(b,x_{\frac{1}{2}})y_{0}b_{0}-(b,x_{\frac{1}{2}}y_{0})b_{0}.$
Proof.
By Lemma 4.11(2),
(7.5) $0=$
$b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})+(2(a,b)-1)(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+4(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})$
$+4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})$
$-(b,y_{0})x_{\frac{1}{2}}b_{\frac{1}{2}}-(b,x_{\frac{1}{2}})y_{0}b_{0}-(b,x_{\frac{1}{2}}y_{0})(a,b)a-(b,x_{\frac{1}{2}}y_{0})b_{0}.$
Computing 22 times (7.3) minus (7.5) we get
$\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}$
$-4\Big((b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-4((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}$
$+(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}+(b,x_{\frac{1}{2}})y_{0}b_{0}+(b,x_{\frac{1}{2}}y_{0})b_{0}=0.\qed$
Proposition 7.4.
  1. (1)
    $\frac{1}{2}(1-(a,b))\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}\Big]=$
    $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)+\Big((b_{\frac{1}{2}}(y_{0}b_{0}))x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0}).$
  2. (2)
    $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})=$
    $\frac{1}{2}(1-(a,b))\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}\Big]-\frac{1}{4}(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}.$
Proof.
We put $y_{0}b_{0}$ in place of $y_{0}$ in (7.4), noticing that by Theorem 1.4, if we denote by $z$ the element after the equality sign in (7.4), then replacing $y_{0}$ by $y_{0}b_{0}$ in $z$ gives
$\textstyle{\frac{1}{2}(1-(a,b))z+2(b,y_{0})\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}+(b,y_{0})(b_{\frac{1}{2}}x_{\frac{1}{2}})b_{0}}$
$\textstyle{-\frac{1}{2}(1-(a,b))(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}-\frac{1}{2}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}-\frac{1}{4}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}.}$
Hence
$\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}+\Big((b_{\frac{1}{2}}(y_{0}b_{0}))x_{\frac{1}{2}}-(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})\Big)_{0}=$
$\textstyle{\frac{1}{2}(1-(a,b))\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}\Big]}$
$+2(b,y_{0})\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}+(b,y_{0})(b_{\frac{1}{2}}x_{\frac{1}{2}})b_{0}$
$\textstyle{-\frac{1}{2}(1-(a,b))(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}-\frac{1}{2}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}-\frac{1}{4}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}}$
$\iff$
$\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}+\Big((b_{\frac{1}{2}}(y_{0}b_{0}))x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})=$
$\textstyle{\frac{1}{2}(1-(a,b))\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}\Big]}$
$\iff$
$\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})$
$\textstyle{\frac{1}{2}(1-(a,b))\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}\Big]-\frac{1}{4}(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}.}\qed$
Lemma 7.5.
  1. (1)
    $\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}=0.$
  2. (2)
    $\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})(y_{0}b_{0}))\Big)_{0}=\\ -\frac{1}{2}(1-(a,b))\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}+\frac{1}{8}(1-(a,b))(b,x_{\frac{1}{2}})y_{0}b_{0}\\ +\frac{1}{16}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}=\\ \frac{1}{2}(1-(a,b))\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}+\frac{1}{16}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}.$
  3. (3)
    If $(a,b)\notin\{\frac{1}{4},1\},$ then
    $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}=-\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}\\ +\frac{1}{4}(b,x_{\frac{1}{2}}y_{0})b_{0}\\ =b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})+\frac{1}{4}(b,x_{\frac{1}{2}}y_{0})b_{0}.$
Proof.
(1) Putting in (7.5) $x_{\frac{1}{2}}b_{0}$ in place of $x_{\frac{1}{2}}$ we get
$0=$
$\textstyle{\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}+(2(a,b)-1)\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}}$
$+2(1-(a,b))\Big((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}+(b,x_{\frac{1}{2}})(y_{0}b_{0})b_{0}$
$\textstyle{-\frac{1}{4}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}-\frac{1}{2}(1-(a,b))(b,x_{\frac{1}{2}})y_{0}b_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}}$
$\textstyle{\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}}$
$\textstyle{+\frac{1}{2}(1-(a,b))(b,x_{\frac{1}{2}})y_{0}b_{0}+\frac{1}{2}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}}$
$\textstyle{-\frac{1}{4}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}-\frac{1}{2}(1-(a,b))(b,x_{\frac{1}{2}})y_{0}b_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}}$
$\textstyle{\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}.}$
(2) Putting in (1) $y_{0}b_{0}$ in place of $y_{0},$ we get
$\textstyle{\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})(y_{0}b_{0}))\Big)_{0}=-\Big((b_{\frac{1}{2}}(y_{0}b_{0}))(x_{\frac{1}{2}}b_{0})\Big)_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})(y_{0}b_{0})b_{0}}$
$\textstyle{=-\frac{1}{2}(1-(a,b))\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}-\frac{1}{16}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}}$
$\textstyle{+\frac{1}{8}(1-(a,b))(b,x_{\frac{1}{2}})y_{0}b_{0}+\frac{1}{8}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}.}$
(3) By Lemma 6.2(1) and part (2), using Theorem 4.9,
$(1-(a,b))\Big(b_{\frac{1}{2}}((y_{\frac{1}{2}}x_{0})b_{0})\Big)_{0}-2\Big(b_{\frac{1}{2}}(((y_{\frac{1}{2}}b_{0})x_{0})b_{0})\Big)_{0}$
$=(1-(a,b))\Big(b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{0}b_{0}))\Big)_{0}-2\Big(b_{\frac{1}{2}}((y_{\frac{1}{2}}b_{0})(b_{0}x_{0}))\Big)_{0}$
$\iff$
$\textstyle{\frac{1}{4}(1-(a,b))(b,y_{\frac{1}{2}}x_{0})b_{0}-\frac{1}{2}(b,(y_{\frac{1}{2}}b_{0})x_{0}))b_{0}}$
$=(1-(a,b))\Big(b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{0}b_{0}))\Big)_{0}+(1-(a,b))\Big((b_{\frac{1}{2}}x_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}$
$\textstyle{-\frac{1}{4}(1-(a,b))(b,y_{\frac{1}{2}})x_{0}b_{0}-\frac{1}{8}(b,x_{0})(b,y_{\frac{1}{2}})b_{0}}$
$\iff$
$\textstyle{\frac{1}{4}(1-(a,b))(b,y_{\frac{1}{2}}x_{0})b_{0}-\frac{1}{8}(b,y_{\frac{1}{2}})(b,x_{0})b_{0}}$
$=(1-(a,b))\Big(b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{0}b_{0}))\Big)_{0}$
$\textstyle{+(1-(a,b))\Big((b_{\frac{1}{2}}x_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}-\frac{1}{4}(1-(a,b))(b,y_{\frac{1}{2}})x_{0}b_{0}}$
$\textstyle{-\frac{1}{8}(b,x_{0})(b,y_{\frac{1}{2}})b_{0},}$
so the first equality in (3) follows from by replacing $y_{\frac{1}{2}}$ with $x_{\frac{1}{2}},$ and $x_{0}$ with $y_{0}.$ The second equality follows from this and (1). ∎
Corollary 7.6.
  1. (1)
    $(2(a,b)-1)\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}\Big]$
    $4b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})+4\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-(b,x_{\frac{1}{2}}y_{0})b_{0}=0.$
  2. (2)
    If $(a,b)\notin\{\frac{1}{4},1\},$ then
    $(2(a,b)-1)\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}\Big]=0$
Proof.
By Proposition 7.4(2), $4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}\quad=\\ 4\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}+2((a,b)-1)\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}\Big].$
Hence by (7.5), and since $\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}=0,$ by Lemma 7.5(1),
$0=$
$\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+(2(a,b)-1)\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+4\Big((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}$
$+4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}$
$-(b,x_{\frac{1}{2}})y_{0}b_{0}-(b,x_{\frac{1}{2}}y_{0})b_{0}$
$(2(a,b)-1)\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}\Big]$
$+4\Big((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}+4\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}$
$-(b,x_{\frac{1}{2}})y_{0}b_{0}-(b,x_{\frac{1}{2}}y_{0})b_{0}$
$(2(a,b)-1)\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}\Big]$
$-4b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})+4\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-(b,x_{\frac{1}{2}}y_{0})b_{0}.$
(2) Follows from (1) and Lemma 7.5(3). ∎
Corollary 7.7.
$m=\frac{1}{2}(1-(a,b))\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}+\frac{1}{16}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}.$
Proof.
By Lemma 7.5(1), $\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}=-\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}.$ Note that by Theorem 4.9, for $z:=-\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0},$ we have
$\textstyle{zb_{0}=\frac{1}{2}(1-(a,b))z-\frac{1}{4}(b,(x_{\frac{1}{2}}b_{0})y_{0})b_{0}+\frac{1}{8}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}}$
$\textstyle{=\frac{1}{2}(1-(a,b))z+\frac{1}{16}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}.}$
But $z=\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0},$ so the colrollary holds. ∎
Proposition 7.8.
  1. (1)
    $(a,b)\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}\Big]$
    $2((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})-2((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}+\frac{1}{2}(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}=0.$
  2. (2)
    Let $z:=\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}.$ If $(a,b)\neq 0,$ then $zb_{0}=\frac{1}{2}(1-(a,b)z.$
Proof.
(1) By (7.3) and Proposition 7.4(2),
$0=$
$\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+(a,b)\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+2\Big((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}$
$\textstyle{-(1-(a,b))\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}\Big]+\frac{1}{2}(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}.}$
$-2((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}$
$(a,b)\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big]$
$\textstyle{+2((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})-2((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}+\frac{1}{2}(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}.}$
(2) By (1), (a,b)z=w,(a,b)z=-w, with $w=2((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})-2((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}+\frac{1}{2}(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}.$ However, by Corollary 7.7, and by Theorem 1.4,
$\textstyle{wb_{0}=\frac{1}{2}(1-(a,b))w+\frac{1}{8}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}-(b_{0},(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}+\frac{1}{8}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}}$
$\textstyle{=\frac{1}{2}(1-(a,b))w.}\qed$
We can now prove Theorem 7.1.
Proof of Theorem 7.1..
By Corollary 7.6(2), we may assume that $(a,b)=\frac{1}{2}.$
Let $z:=\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}.$ By Proposition 7.8(2), $zb_{0}=\frac{1}{2}(1-(a,b))z.$ By (7.1),
${z=2zb_{0}+2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))-2b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})-\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}.}$
But by Lemma 7.5(3), $2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))-2b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})-\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}=0.$ Hence, by Proposition 7.8(2), $z=2zb_{0}=(1-(a,b))z.$ It follows that $(a,b)z=0,$ so $z=0.$
8. The case $(a,b)=1$
In this section we assume that $a,b$ are axes and that $(a,b)=1.$
Theorem 8.1.
  1. (1)
    If $b_{0}=0,$ then $L_{a},L_{b}]$ is a derivation.
  2. (2)
    Assume ${\rm char}(\mathbb{F})\neq 3.$ Then
    • (i)
      $(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}=$
      $\frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
    • (ii)
      $b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})+(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}=0.$
In particular, $L_{a},L_{b}]$ is a derivation.
Proof. (1) Suppose $b_{0}=0.$ Then, by Lemma 4.3(3), $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}=(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-\Big(x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0})\Big)_{0}.$ By Lemma 4.13(2), $(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}=\frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}},$ because $(b,x_{\frac{1}{2}}y_{\frac{1}{2}})=(a,x_{\frac{1}{2}}y_{\frac{1}{2}}).$ So (1) follows from Theorem 5.1, and Theorem 3.7.
(2) Note that since $(a,b)=1,$ theorem 4.9 implies
(8.1) $(1)\ (x_{\frac{1}{2}}b_{0})b_{0}=0.$
$(2)\ \textstyle{(x_{0}b_{0})b_{0}=\frac{1}{2}(b,x_{0})b_{0}.}$
$(3)\ \textstyle{b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})=\frac{1}{2}(b,x_{0})a+\frac{1}{2}x_{0}b_{0}.}$
$(4)\ \textstyle{b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{\frac{1}{2}})=-\frac{1}{2}x_{\frac{1}{2}}b_{0}+\frac{1}{2}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}\text{ and}}$
$\quad\ \ \textstyle{b_{\frac{1}{2}}\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}=-\frac{1}{2}x_{\frac{1}{2}}b_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}.}$
$(5)\ \textstyle{b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})=(b_{\frac{1}{2}}x_{\frac{1}{2}})b_{0}=\frac{1}{4}(b,x_{\frac{1}{2}})b_{0}.}$
$(6)\ \textstyle{b_{\frac{1}{2}}(x_{0}b_{0})=\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}.}$
$(7)\ (b_{0},(x_{\frac{1}{2}}b_{0})y)=0=(b_{\frac{1}{2}},yb_{0}),\ \forall y\in A.$
[…]
Lemma 8.2.
  1. (1)
    $(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})=\frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}+\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})).$
  2. (2)
    $4(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})+4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})=\\ (b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}+4(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-4b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))-4b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})).$
  3. (3)
    $((b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0}))b_{0}=0.$
  4. (4)
    $((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}=\frac{1}{2}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]-\frac{1}{4}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
  5. (5)
    $4(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})+4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})\quad=\\ (b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}-\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))+2b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})).$
  6. (6)
    $6\Big[b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))+b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))\Big]=\frac{9}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
Proof.
(1) By Lemma 4.13(2),
(8.2) $(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}+4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})+4(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})\quad=$
$\textstyle{\frac{1}{2}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}\Big]+(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
$+(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
Replacing $x_{\frac{1}{2}}$ by $x_{\frac{1}{2}}b_{0}$ in (8.2) using (8.1), we get
$\textstyle{\frac{1}{4}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}+(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})+0+0\quad=}$
$\textstyle{\frac{1}{2}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+0\Big]+0+0}$
$\textstyle{+(a,(x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})b_{\frac{1}{2}}-b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}}),}$
so (1) holds.
(2) This follows by interchanging $x_{\frac{1}{2}}$ and $y_{\frac{1}{2}}$ in (1) and adding.
(3) Multiplying (1) by $b_{0}$ using (8.1) we get $((b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0}))b_{0}=0-0+0-(b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})))b_{0}=\frac{1}{4}(b_{0},(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))b_{0}=0.$
(4) Multiplying (8.2) by $b_{0}$ using (3) and (8.1)(6) we get
$((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}+((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+0+0\quad=$
$\textstyle{\frac{1}{2}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]+0+0}$
$\textstyle{+0-\frac{1}{4}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
(5) By Lemma 4.5(3b) we have
$+4(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})+4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})\quad=$
$2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))+2b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))$
$+2((b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0}+2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}.$
Using (4) we get
$+4(b_{\frac{1}{2}}y_{\frac{1}{2}})(x_{\frac{1}{2}}b_{0})+4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})\quad=$
$2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))+2b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))$
$\textstyle{+(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}-\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
(6) By (2) and (5),
$4(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-4\Big[b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))+b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))\Big]=$
$\textstyle{-\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+2\Big[b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))+b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))\Big],}$
so (6) holds. ∎
Proof of Theorem 8.1(2i).
By (8.2) and Lemma 8.2(2),
$b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}=\\ \frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-3(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})+4\Big[b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))+b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))\Big],$ so (2i) follows from Lemma 8.2(6), since ${\rm char}(\mathbb{F})\neq 3.$
For the proof of part (2ii), we need the following three lemmas.
Lemma 8.3.
  1. (1)
    $\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}+2(x_{\frac{1}{2}}b_{\frac{1}{2}})(y_{0}b_{0})=\\ \frac{3}{4}(b,y_{0})\Big(x_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{0}b_{0}-(x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0}))b_{0}+\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}.$
  2. (2)
    $\Big((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}+((x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})b_{0}-\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}=0.$
  3. (3)
    $(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})b_{0}=\frac{1}{16}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}.$
  4. (4)
    $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}+\frac{1}{4}(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})=0.$
  5. (5)
    $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+\\ 2\Big((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}+\frac{1}{2}(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}-2((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}=0.$
  6. (6)
    $\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}=0.$
  7. (7)
    • (i)
      $\Big[b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}\Big]b_{0}=\\ \frac{1}{4}(b,x_{\frac{1}{2}}y_{0})b_{0}+((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}-((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}=0.$
    • (ii)
      $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}=\\ 2\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-2\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}-\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}.$
    • (iii)
      $((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}-\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}=-\frac{3}{4}(b,x_{\frac{1}{2}}y_{0})b_{0}+((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}.$
Proof.
By Lemma 4.13(1),
(8.3) $\Big(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0})\Big)_{0}+\Big(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}+4\Big((x_{\frac{1}{2}}b_{\frac{1}{2}})(y_{\frac{1}{2}}b_{\frac{1}{2}})\Big)_{0}$
$+4\Big((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}\quad=$
$(b,y_{\frac{1}{2}})\Big(x_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}+(b,x_{\frac{1}{2}})\Big(y_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}-(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}+(b,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}.$
(1) Replacing in (8.3) $y_{\frac{1}{2}}$ by $b_{\frac{1}{2}}y_{0}$ we get
$\textstyle{\frac{1}{4}(b,y_{0})\Big(x_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}+2(x_{\frac{1}{2}}b_{\frac{1}{2}})(y_{0}b_{0})+\frac{1}{4}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}=}$
$\textstyle{(b,y_{0})\Big(x_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{0}b_{0}-(x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0}))b_{0}+\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}+\frac{1}{4}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}.}$
(2) Replacing $x_{\frac{1}{2}}$ with $x_{\frac{1}{2}}b_{0}$ in (8.3), using (8.1) we get
$\textstyle{\Big((x_{\frac{1}{2}}b_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}+0+\frac{1}{4}(b,x_{\frac{1}{2}})(b,y_{\frac{1}{2}})b_{0}+0\quad=}$
$\textstyle{\frac{1}{4}(b,x_{\frac{1}{2}})(b,y_{\frac{1}{2}})b_{0}+0-((x_{\frac{1}{2}}b_{0})y_{\frac{1}{2}})b_{0}+\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{0}.}$
(3) Replacing in (1) $x_{\frac{1}{2}}$ with $x_{\frac{1}{2}}b_{0}$ we get
$\textstyle{0+\frac{1}{4}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}=}$
$\textstyle{\frac{3}{16}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}+0-((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0}))b_{0}+\frac{1}{8}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}.}$
(4&5) By Lemma 4.3(3),
(*) $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+2\Big((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}$
$+2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-2\Big((b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0})\Big)_{0}-2((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}=0.$
Replacing $y_{0}$ with $y_{0}b_{0}$, and using (8.1), we get
$\textstyle{\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}+\frac{1}{4}(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})+\frac{1}{8}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}}$
$\textstyle{+\frac{1}{4}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}-\frac{1}{4}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}-\frac{1}{8}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}=0,}$
so (4) holds; going back to (*) we get (5).
(6) Replacing $x_{\frac{1}{2}}$ with $x_{\frac{1}{2}}b_{0}$ in (5), we get
$\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}-(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))y_{0}+0$
$\textstyle{+\frac{1}{2}(b,y_{0})\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}-2((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0}))b_{0}=0.}$
So (6) follows from (3) and Theorem 4.9(5).
(7) (i) Multiplying (5) by $b_{0}$ using (8.1) we get
$\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}+(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}\Big]b_{0}$
$\textstyle{+\frac{1}{8}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}+\frac{1}{8}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}-(b_{0},(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}=0.}$
Since $(b_{0},(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})=\frac{1}{4}(b,x_{\frac{1}{2}})(b,y_{0})b_{0},$ and $b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})b_{0}=\frac{1}{4}(b,x_{\frac{1}{2}}y_{0})b_{0},$ (7i) holds.
(ii) This follows from (7.1) and (i).
(iii) This follows from (i).
Lemma 8.4.
$2(b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})+2(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-2((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}$
$-2b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})=0.$
Proof.
This follows from equation (7.2). ∎
Lemma 8.5.
  1. (1)
    $(x_{0}b_{0})(y_{0}b_{0})=(x_{0}(y_{0}b_{0}))b_{0}.$
  2. (2)
    $(x_{0}y_{0})b_{0}+2\Big((b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}=y_{0}(x_{0}b_{0})+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{0})x_{0})\Big)_{0}.$
  3. (3)
    $4\Big((b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}=y_{0}(x_{0}b_{0})+x_{0}(y_{0}b_{0})-(x_{0}y_{0})b_{0}.$
  4. (4)
    • (i)
      $2\Big((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}+(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}=0.$
    • (ii)
      $2(b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})+2b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})-\frac{1}{2}(b,x_{\frac{1}{2}})y_{0}b_{0}=0.$
  5. (5)
    If ${\rm char}(\mathbb{F})\neq 3,$ then
    • (i)
      $\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}+(x_{\frac{1}{2}}b_{\frac{1}{2}})(y_{0}b_{0})-\frac{1}{4}(b,x_{\frac{1}{2}}y_{0})b_{0}\\ =\frac{1}{4}(b,y_{0})\Big(x_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}.$
    • (ii)
      $\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}+\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-\frac{1}{4}(b,x_{\frac{1}{2}}y_{0})b_{0}\\ =\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}.$
    • (iii)
      $\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}+\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-\frac{1}{4}(b,x_{\frac{1}{2}}y_{0})b_{0}=0.$
Proof.
(1&2): By Lemma 4.2(9),
$(x_{0}y_{0})b_{0}+2(x_{0}b_{0})(y_{0}b_{0})+2\Big((b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}$
$=y_{0}(x_{0}b_{0})+2(x_{0}(y_{0}b_{0}))b_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{0})x_{0})\Big)_{0}.$
Replacing $y_{0}$ with $y_{0}b_{0}$ using (8.1), we get
$\textstyle{(x_{0}(y_{0}b_{0}))b_{0}+\frac{1}{2}(b,x_{0})(b,y_{0})b_{0}+\frac{1}{4}(b,y_{0})x_{0}b_{0}}$
$\textstyle{=(y_{0}b_{0})(x_{0}b_{0})+\frac{1}{2}(b,x_{0})(b,y_{0})b_{0}+\frac{1}{4}(b,y_{0})x_{0}b_{0}.}$
This shows (1), and then (2) follows.
(3) Interchanging $x_{0},y_{0}$ in (2) and adding, and then using Theorem 5.1 and (8.1)(5), we get
$2(x_{0}y_{0})b_{0}+4\Big((b_{\frac{1}{2}}x_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}$
$=y_{0}(x_{0}b_{0})+x_{0}(y_{0}b_{0})+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}y_{0})x_{0})\Big)_{0}+2\Big(b_{\frac{1}{2}}((b_{\frac{1}{2}}x_{0})y_{0})\Big)_{0}$
$y_{0}(x_{0}b_{0})+x_{0}(y_{0}b_{0})+(x_{0}y_{0})b_{0}.$
(4) (i) Putting $\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}$ in place of $x_{0}$ in (3) we get using (8.1),
$\textstyle{4\Big(\Big[-\frac{1}{2}x_{\frac{1}{2}}b_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}\Big](b_{\frac{1}{2}}y_{0})\Big)_{0}}$
$\textstyle{=\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}+(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}}$
$\iff$
$\textstyle{-2\Big((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{0}b_{0}}$
$\textstyle{=\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}+(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}.}$
(ii) Computing 22 times the equality in (i) minus the equality in Lemma 8.4 yields (ii).
(5) (i) By Lemma 8.3 (1) and (7)(iii),
$\textstyle{\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}+2(x_{\frac{1}{2}}b_{\frac{1}{2}})(y_{0}b_{0})-\frac{3}{4}(b,x_{\frac{1}{2}}y_{0})b_{0}+((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}=}$
$\textstyle{\frac{3}{4}(b,y_{0})\Big(x_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{0}b_{0}.}$
Adding (4i) we get
$\textstyle{3\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}+3(x_{\frac{1}{2}}b_{\frac{1}{2}})(y_{0}b_{0})-\frac{3}{4}(b,x_{\frac{1}{2}}y_{0})b_{0}}$
$\textstyle{=\frac{3}{4}(b,y_{0})\Big(x_{\frac{1}{2}}b_{\frac{1}{2}}\Big)_{0}+\frac{3}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}.}$
(ii) By Lemma 8.3(4), $(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-\frac{1}{4}(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}=\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0},$ so (ii) follows from (i).
(iii) By Lemma 8.3(6), $\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}=-\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0},$ so (iii) follows from (ii).∎
Proof of Theorem 8.1(2ii).
The proof of this part is immediate from Lemma 8.3(7ii) and Lemma 8.5(5iii). ∎